### v1

# Regularized Extragradient Methods for Solving Equilibrium Problems on Hadamard Manifolds

Abstract
Employing two distinct types of regularization terms, we propose two regularized extragradient methods for solving equilibrium problems on Hadamard manifolds. The sequences generated by these extragradient algorithms converge to a solution of the equilibrium problem without requiring the Lipschitz continuity of the bifunction or imposing additional conditions on the parameters. We establish convergence results for both algorithms and derive global error bounds along with $R$-linear convergence rates in cases where the bifunction is strongly pseudomonotone. Finally, we present numerical experiments to demonstrate the effectiveness of our methods.
Keywords: Hadamard manifold, Equilibrium Problem, Busemann function, Extragradient method, Linear convergence
1 Introduction
The equilibrium problem (EP) has been extensively studied and remains a highly active area of research. One of the key motivations behind its investigation is its broad applicability, since numerous mathematical problems can be formulated as equilibrium problems. These include optimization problems, Nash equilibrium problems, complementarity problems, fixed point problems, and variational inequality problems. A comprehensive discussion of EP and its developments can be found in the works of Blum and Oettli [1], Binachi and Schaible [2], as well as in references therein.
The study of the proximal point method on Hadamard manifolds has gained considerable attention, with various researchers exploring its application to specific cases of the equilibrium problem (EP); see, for instance, the works of Ferreira and Oliveira [3], Li et al. [4], and Tang et al. [5]. More recently, there has been a natural shift toward extending classical Euclidean concepts and methodologies to the Riemannian setting. Notable contributions in this direction include those by Ledyaev and Zhu [6], Ferreira et al. [7], Li et al. [8, 9], among others. A significant advantage of adapting these techniques to Riemannian manifolds is the ability to reformulate certain nonconvex problems as convex ones by selecting an appropriate Riemannian metric. Relevant discussions on this topic can be found in the papers by Rapcsák [10], Cruz Neto et al. [11], Bento and Melo [12], and by Colao et al. [13].
Note that Hadamard manifolds generally do not have a linear structure, which indicates that properties, techniques as well as algorithms in linear spaces cannot be used in Hadamard manifolds. Therefore, it is valuable and interesting to extend algorithms for solving equilibrium problems from linear spaces to Hadamard manifolds.
Let $\mathbb{M}$ be an Hadamard manifold. Given $C\subset\mathbb{M}$, a nonempty, closed and convex set, and a bifunction $F:C\times C\to\mathbb{R}$ satisfying the property $F(x,x)=0$ for all $x\in C$, the equilibrium problems in the Riemanninan context (denoted by $\operatorname*{EP}(F,C)$) is formulated as follows:
Find $\text{Find }x^{*}\in C\text{ such that }F(x^{*},y)\geq 0\text{ for all }y\in C.$ (1.1)
In recent years various extragradient-based methods have been developed in order to solve equilibrium problem (1.1) on Hadamard manifolds. Colao et al. [13] established the existence of solutions to such problems under suitable conditions on the bifunction. Building upon their work, Cruz Neto et al. [14] proposed an extragradient algorithm with a fixed step size and proved its global convergence. Inspired by the works of Cruz Neto et al. [14] and Hieu [15, 16], Khammahawong et al. [17] introduced two extragradient methods employing decreasing and non-summable step sizes, extending Hieu’s results from Hilbert space to the Hadamard setting. For strongly pseudomonotone bifunctions, they demonstrated convergence of the generated sequences under mild assumptions.
Further extending this line of research, Ansari et al. [18] proposed two adaptive explicit extragradient algorithms for solving EP (1.1) on Hadamard manifolds. Their methods utilize a simple, information-based adaptive step size rule, eliminating the need for prior knowledge of the Lipschitz constant. Convergence was established under pseudomonotonicity, with linear convergence shown for strongly pseudomonotone bifunctions.
It is worth noting that the methods in [14, 17, 18] require solving a strongly convex optimization problem twice per iteration. To address this computational burden, Chen et al. [19] introduced a modified golden ratio algorithm that reduces this requirement to once per iteration. Numerical experiments have shown that their method outperforms those of [14, 17] in terms of efficiency and computational cost.
In 2022, Babu et al. [20] developed several extragradient algorithms incorporating an Armijo-type step size rule for solving equilibrium problems on Hadamard manifolds. More recently, Oyewole and Reich [21] introduced an inertial subgradient extragradient method with a self-adaptive step size. However, these methods tend to be computationally intensive, as the associated strongly convex optimization problem may need to be solved multiple times at each iteration.
In 2022, Bento et al. [22] introduced a regularization of the proximal point algorithm in the context of Hadamard manifolds using a new regularization term which has a clear interpretation in a recent variational rationality approach to human behavior. Later, Bento et al. [23] introduced a new resolvent for the bifunction in terms of Busemann functions. An advantage of this new proposal is that, in addition to being a natural extension of its counterpart in the linear setting introduced by Combettes and Hirstoaga [24], the new term that performs the regularization is a convex function in general Hadamard manifolds. Bento et al. [25] have recently considered the resolvent via Busemann functions introduced by Bento et al. [23] and presented a proximal point method for solving equilibrium problems on Hadamard manifolds.
Motivated by the works of Bento et al. [22], Bento et al. [23] and Bento et al. [25], we propose two regularized extragradient methods incorporating distinct regularization terms for solving equilibrium problems on Hadamard manifolds. The proposed algorithms operate without any step size restrictions, do not require a line search procedure, and do not assume Lipschitz continuity of the bifunction, thus making them suitable for solving equilibrium problems involving monotone bifunctions.
2 Preliminaries
For the basic definitions and results regarding the geometry of Riemannian manifolds, we refer to [26, 27, 28]. Let $\mathbb{M}$ be an $m$-dimensional differentiable manifold and let $x\in\mathbb{M}$. The set of all tangent vectors at the point $x$, denoted by $T_{x}\mathbb{M}$, is called the tangent space of $\mathbb{M}$ at $x$. It forms a real vector space of dimension $m$. The tangent bundle of $\mathbb{M}$ is defined as the disjoint union $T\mathbb{M}=\bigcup_{x\in\mathbb{M}}T_{x}\mathbb{M}$.
[…]
A single-valued vector field on a differentiable manifold $\mathbb{M}$ is a smooth mapping $A\colon\mathbb{M}\to T\mathbb{M}$ such that, for each $x\in\mathbb{M}$, there exists a tangent vector $A(x)\in T_{x}\mathbb{M}$. Let $\triangle$ be the Levi-Civita connection associated with the Riemannian manifold $M$. A vector field $A$ along $\gamma$ is said to be parallel if $\triangle_{\gamma^{\prime}}A=0$. If $\gamma^{\prime}$ is parallel along $\gamma$, that is, $\triangle_{\gamma^{\prime}}\gamma^{\prime}=0$, then $\gamma$ is said to be geodesic. In this case, $\|\gamma^{\prime}\|$ is a constant. Moreover, if $\|\gamma^{\prime}\|=1$, then $\gamma$ is called a normalized geodesic. A geodesic joining $x$ to $y$ in the Riemannian manifold $\mathbb{M}$ is said to be a minimal geodesic if its length is equal to $d(x,y)$. We denote the geodesic joining $x$ and $y$ by $\gamma(x,y;\cdot)$, that is, $\gamma(x,y;\cdot):[a,b]\to\mathbb{M}$ is such that, for $a,b\in\mathbb{R}$, we have $\gamma(x,y;a)=x$ and $\gamma(x,y;b)=y$.
The Riemannian metric induces a mapping $f\mapsto\operatorname*{grad}f$ that associates to each differentiable function $f\colon\mathbb{M}\to\mathbb{R}$ a unique vector field $\operatorname*{grad}f\colon\mathbb{M}\to T\mathbb{M}$, called the gradient of $f$, defined by $\langle\operatorname*{grad}f(x),X(x)\rangle=df_{x}(X(x))$ for all vector fields $X\colon\mathbb{M}\to T\mathbb{M}$ and all $x\in\mathbb{M}$. Here, $df_{x}$ denotes the differential of $f$ at the point $x\in\mathbb{M}$.
A Riemannian manifold $\mathbb{M}$ is said to be complete if for any $x\in\mathbb{M}$, each geodesic emanating from $x$ is defined for all $t\in\mathbb{R}$. By the Hopf-Rinow theorem [27, Theorem 1.1], if $\mathbb{M}$ is complete, then each pair of points in $\mathbb{M}$ can be joined by a minimal geodesic. Moreover, $(\mathbb{M},d)$ is a complete metric space. If $\mathbb{M}$ is a complete Riemannian manifold, then the exponential map $\exp_{x}\colon T_{x}\mathbb{M}\to\mathbb{M}$ at $x\in\mathbb{M}$ is defined by $\exp_{x}v=\gamma_{x,v}(1)\text{ for all }v\in T_{x}\mathbb{M}$, where $\gamma_{x,v}\colon\mathbb{R}\to\mathbb{M}$ is the unique geodesic starting from $x$ with velocity $v$, that is, $\gamma_{x,v}(0)=x$ and $\gamma^{\prime}_{x,v}(0)=v$. It is known that $\exp_{x}(tv)=\gamma_{x,v}(t)$ for each real number $t$ and $\exp_{x}(\mathbf{0})=\gamma_{x,\mathbf{0}}(0)=x$, where $\mathbf{0}$ is the zero tangent vector.
Definition 2.1 ([29]).
A complete simply connected Riemannian manifold of nonpositive sectional curvature is called an Hadamard manifold.
Definition 2.2 ([29]).
A nonempty subset $C$ of an Hadamard manifold $\mathbb{M}$ is said to be geodesically convex if, for any two points $x$ and $y$ in $C$, the geodesic joining $x$ to $y$ is contained in $C$, that is, if $\gamma(x,y;\cdot)\colon[a,b]\to\mathbb{M}$ is the geodesic such that $x=\gamma(x,y;a)$ and $y=\gamma(x,y;b)$, then $\gamma((1-t)a+tb)\in C$ for all $t\in[0,1]$.
Definition 2.3 ([29]).
Let $\mathbb{M}$ be an Hadamard manifold. A function $f\colon\mathbb{M}\to\mathbb{R}$ is said to be geodesically convex if, for any geodesic $\gamma(x,y;\cdot)\colon[a,b]\to\mathbb{M}$, the composite function $f\circ\gamma(x,y;\cdot)\colon[a,b]\to\mathbb{R}$ is convex, that is,
$f\circ\gamma(x,y;(1-t)a+tb)\leq(1-t)(f\circ\gamma(x,y;a))+t(f\circ\gamma(x,y;b))$
for any $a,b\in\mathbb{R}$, $x,y\in\mathbb{M}$ and $t\in[a,b]$.
Lemma 2.4 ([29]).
Let $\{x_{n}\}$ be a sequence in an Hadamard manifold $\mathbb{M}$ such that $x_{n}\to x\in\mathbb{M}$. Then, for any $y\in\mathbb{M}$, we have
$\exp_{x_{n}}^{-1}y\to\exp_{x}^{-1}y\mbox{ and }\exp_{y}^{-1}x_{n}\to\exp_{y}^{-1}x.$
Remark 2.5 ([29]).
Let $\mathbb{M}$ be an Hadamard manifold and let $x,y,z\in\mathbb{M}$. Then
$d^{2}(x,y)+d^{2}(y,z)-d^{2}(z,x)\leq 2\langle\exp_{y}^{-1}x,\exp_{y}^{-1}z\rangle.$
Definition 2.6 ([28]).
Let $\mathbb{M}$ be an Hadamard manifold and let $C$ be a nonempty, geodesically convex subset of $\mathbb{M}$. Let $g\colon C\to\mathbb{R}$ be a real-valued function. Then the subdifferential $\partial g\colon\mathbb{M}\to T\mathbb{M}$ of $g$ at $x$ is defined by
$\partial g(x)\coloneqq\left\{v\in T_{x}\mathbb{M}:\langle v,\exp_{x}^{-1}y\rangle\leq g(y)-g(x)\text{ for all }y\in C\right\}.$
Definition 2.7 ([30]).
Let $C$ be a nonempty subset of an Hadamard manifold $\mathbb{M}$ and let $f\colon C\to(-\infty,\infty]$ be a proper, geodesically convex and lower semicontinuous function. The proximal operator of $f$ is defined by
$\operatorname*{prox}_{\lambda f}(x)\coloneq\operatorname*{argmin}_{y\in C}\left(f(y)+\frac{1}{2\lambda}d^{2}(y,x)\right)\text{ for all }x\in C\text{ andfor any }\lambda\in(0,\infty).$
Lemma 2.8 ([3], Theorem 5.1).
Let $\mathbb{M}$ be an Hadamard manifold and let $f:\mathbb{M}\to\mathbb{R}$ be a geodesically convex function. Let $\{x_{n}\}$ be the sequence generated by the proximal point algorithm, that is,
$x_{n+1}=\operatorname*{argmin}_{y\in\mathbb{M}}\left\{f(y)+\frac{1}{2\lambda_{n}}d^{2}(x_{n},y)\right\}\text{ for all }n\in\mathbb{N}_{0}$
with the initial point $x_{0}\in\mathbb{M}$ and $\lambda_{n}\in(0,\infty)$. Then the sequence $\{x_{n}\}$ is well defined and characterized by $\frac{1}{\lambda_{n}}\exp_{x_{n+1}}^{-1}x_{n}\in\partial f(x_{n+1}).$
Remark 2.9.
In view of [3, Lemma 4.2], $\operatorname*{prox}_{\lambda f}(x)$ is single-valued. By the definition of $\partial f(x_{n+1})$ and Lemma 2.8, we have
$\frac{1}{\lambda_{n}}\langle\exp_{x_{n+1}}^{-1}x_{n},\exp_{x_{n+1}}^{-1}x\rangle\leq f(x)-f(x_{n+1})\text{ for all }x\in\mathbb{M}.$
Remark 2.10.
Note that $\operatorname*{grad}d(y,x_{n})=-\frac{\exp_{y}^{-1}x_{n}}{d(y,x_{n})}$ provided $y\neq x_{n}$.
Definition 2.11 ([22, 31]).
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$. A bifunction $F:C\times C\to\mathbb{R}$ is said to be
  1. (i)
    monotone if $F(x,y)+F(y,x)\leq 0$ for all $x,y\in C$;
  2. (ii)
    pseudomonotone if $F(x,y)\geq 0$ implies $F(y,x)\leq 0$ for all $x,y\in C$;
  3. (iii)
    $\alpha$-strongly monotone if there exists a constant $\alpha>0$ such that $F(x,y)+F(y,x)\leq-\alpha d^{2}(x,y)$ for all $x,y\in C$;
  4. (iv)
    $\beta$-strongly pseudomonotone if there exists a constant $\beta>0$ such that $F(x,y)\geq 0$ implies $F(y,x)\leq-\beta d^{2}(x,y)$ for all $x,y\in C$.
Definition 2.12 ([32], Busemann Function).
Let $\mathbb{M}$ be an Hadamard manifold and let $\gamma_{z,x}:[0,\infty)\to\mathbb{M}$ be a geodesic ray parametrized by arc length starting from $z$ and passing through $x$. The Busemann function associated with $\gamma_{z,x}$ is defined by
$b_{\gamma_{z,x}}(y)\coloneqq\lim_{t\to+\infty}\left[d(y,\gamma_{z,x}(t))-t\right]\text{ for all }y\in\mathbb{M}.$
Definition 2.13 ([23]).
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction. The Busemann resolvent $J_{\lambda}^{F}:C\to\mathbb{M}$ associated with $F$ is given by
$J_{\lambda}^{F}(x)\coloneqq\{z\in C:\lambda F(z,y)+d(z,x)b_{\gamma_{z,x}}(y)\geq 0\text{ for all }y\in C\}$ (2.1)
for all $x\in C$ and for any $\lambda\in(0,\infty)$.
As noted in [33, page 273], in the Euclidean setting we have
$d(x,z)b_{\gamma_{x,z}}(y)=\langle x-z,y-x\rangle.$
This shows that the new term, which regularizes the resolvent, is a geodesically convex function. It represents a natural extension of its counterpart introduced in [24]. From [34, Lemma 3.4] $b_{\gamma_{x,z}}(\cdot)$ is Lipschitz continuous with Lipschitz constant 1 and geodesically convex function with $\|\operatorname{grad}b_{\gamma_{x,z}}(\cdot)\|=1$ Moreover, it is known that any linear affine function in the Poincaré plane is constant; see [28, Theorem 2.2, page 299]. Furthermore, in accordance with [35, Theorem 1], this result can be extended to more general Hadamard manifolds with negative sectional curvature. Therefore, Busemann functions seem well-suited to assume the role typically assigned to affine functions in linear settings. As a result, the resolvent in (2.1), aligned with its linear counterpart, proves highly suitable in the Riemannian context.
The Busemann functions, initially introduced by Herbert Busemann to formulate the parallel axiom in a specific class of metric spaces [36], have found applications beyond their original purpose; see, for example, [32, 37, 38, 39] and references therein. These functions have more recently been investigated in the context of optimization [23, 35]. This exploration has spurred ambitious attempts to extend the use of Busemann functions by considering theories and methods of continuous optimization and non-constant affine functions in linear settings.
Assumption 2.14.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction. We consider the following assumptions:
  1. (A1)
    $F$ is monotone and $F(x,x)=0$ for all $x\in C$;
  2. (A2)
    $F(x,\cdot)$ is geodesically convex and sudifferntiable on $C$ for each $x\in C$;
  3. (A3)
    $F(\cdot,y)$ is upper semicontinuous on $C$ for each $y\in C$, that is, $\limsup_{n\to\infty}F(x_{n},y)\leq F(x,y)$ for each $y\in C$ and each $\{x_{n}\}\subset C$ with $x_{n}\to x$;
  4. (A4)
    Given a fixed $z_{0}\in\mathbb{M}$, consider a sequence $\{z_{n}\}\subset C$ such that $\{d(z_{n},z_{0})\}$ tends to $\infty$ as $n$ goes to $\infty$. Then there exist $x^{*}\in C$ and $n_{0}\in\mathbb{N}$ such that
    $F(z_{n},x^{*})\leq 0\text{ for all }n\geq n_{0}.$
Let $\lambda\in(0,\infty)$ and $x\in C$ and let $F:C\times C\to\mathbb{R}$ be a bifunction. For each $x\in C$, consider the set
$K_{x}\coloneqq\{z\in C:\lambda F(z,y)+(d^{2}(y,z)-d^{2}(x,z))\geq 0\}.$
Then, using [23, Theorem 4.1], [25, Theorem 1.1], and [22, Corollary 1], we have the following result:
Theorem 2.15.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction such that Assumptions (A1), (A2), (A3) and (A4) hold. Let $\lambda\in(0,\infty)$. Then
$\operatorname*{EP}(F,C)\neq\emptyset$,
$J_{\lambda}^{F}\neq\emptyset$ for all $x\in\mathbb{M}$,
$J_{\lambda}^{F}$ is single-valued,
the set $K_{x}\neq\emptyset$ for each $x\in C$,
the set $K_{x}$ is a singleton for each $x\in C$.
Theorem 2.16 ([31], Theorem 4.1).
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a $\beta$-strongly monotone bifunction such that Assumptions (A2) and (A3) hold. Then equilibrium problem (1.1) has a unique solution.
Definition 2.17.
Let $C$ be a nonempty, subset of a complete metric space $X$. A sequence $\{x_{n}\}\subset X$ is called Fejér monotone with respect to $C$ if
$d(x_{n+1},y)\leq d(x_{n},y)\text{ for all }y\in C\text{ and for all }n\in\mathbb{N}_{0}.$
Lemma 2.18.
Let $C$ be a nonempty subset of a complete metric space $X$. If $\{x_{n}\}\subset X$ is Fejér monotone with respect to $C$, then $\{x_{n}\}$ is bounded. In addition, if $\{x_{n}\}$ has a cluster point $x$ in $C$, then $\{x_{n}\}$ converges to $x$.
3 Regularized extragradient methods for equilibrium problems
In this section we present two regularized extragradient methods for solving the equilibrium problems (1.1) on Hadamard manifolds.
Algorithm 3.1.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction.
Initialization: Choose $x_{0}\in C$ and $\lambda_{0}\in(0,\infty)$.
Iterative Steps: Given the current iterate $x_{n}$, compute the next iterate as follows:
$\begin{cases}y_{n}=\left\{x\in C:\lambda_{n}F(x,y)+d(x,x_{n})b_{\gamma_{x,x_{n}}}(y)\geq 0\text{ for all }y\in C\right\}=J_{\lambda_{n}}^{F}(x_{n}),\\ x_{n+1}=\operatorname*{argmin}_{y\in C}\left\{F(y_{n},y)+\frac{1}{2\lambda_{n}}d^{2}(x_{n},y)\right\}=\operatorname*{prox}_{\lambda_{n}F(y_{n},\cdot)(x_{n})}\text{ for all }n\in\mathbb{N}_{0},\end{cases}$ (3.1)
where $\{\lambda_{n}\}$ is a sequence in $(0,\infty)$ such that $\limsup_{n\to\infty}\frac{1}{\lambda_{n}}\geq\frac{1}{\lambda}$ for some $\lambda\in(0,\infty)$.
Algorithm 3.2.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction.
Initialization: Choose $x_{0}\in C$ and $\lambda_{0}\in(0,\infty)$.
Iterative Steps: Given the current iterate $x_{n}$, compute the next iterate as follows:
$\begin{cases}y_{n}=\left\{x\in C:\lambda_{n}F(x,y)+(d^{2}(y,x_{n})-d^{2}(x,x_{n}))\geq 0\text{ for all }y\in C\right\},\\ x_{n+1}=\operatorname*{argmin}_{y\in C}\left\{F(y_{n},y)+\frac{1}{2\lambda_{n}}d^{2}(x_{n},y)\right\}=\operatorname*{prox}_{\lambda_{n}F(y_{n},\cdot)(x_{n})}\text{ for all }n\in\mathbb{N}_{0},\end{cases}$ (3.2)
where $\{\lambda_{n}\}$ is a sequence in $(0,\infty)$ such that $\limsup_{n\to\infty}\frac{1}{\lambda_{n}}\geq\frac{1}{\lambda}$ for some $\lambda\in(0,\infty)$.
Lemma 3.3.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction satisfying the assumptions (A1)-(A4). Let $\{x_{n}\}$ and $\{y_{n}\}$ be sequences generated by Algorithm 3.1. Then, for each $x^{*}\in\operatorname*{EP}(F,C)$, we have
$d^{2}(x_{n+1},x^{*})\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n}).$ (3.3)
Furthermore, $\{x_{n}\}$ is Fejér monotone with respect to $\operatorname*{EP}(F,C)$ and the sequences $\{x_{n}\}$ and $\{y_{n}\}$ are bounded.
Proof.
Using the definition of $x_{n+1}$ in Algorithm 3.1 and Remark 2.9, we have
$\lambda_{n}(F(y_{n},y)-F(y_{n},x_{n+1}))\geq\langle\exp^{-1}_{x_{n+1}}x_{n},\exp^{-1}_{x_{n+1}}y\rangle\text{ for all }y\in C.$ (3.4)
Using the definition of $y_{n}$ and $J_{\lambda_{n}}^{F}(x_{n})$, we have
$0\leq\lambda_{n}F(y_{n},y)+d(y_{n},x_{n})b_{\gamma_{y_{n},x_{n}}}(y)\text{ forall }y\in C.$ (3.5)
Since $b_{\gamma_{y_{n},x_{n}}}(y)\leq d(y,x_{n})-d(x_{n},y_{n})$ for all $y\in C$, by (3.5), we have
$0\leq\lambda_{n}F(y_{n},y)+d(y_{n},x_{n})[d(y,x_{n})-d(x_{n},y_{n})]\text{ forall }y\in C.$ (3.6)
Let $G_{n}(y_{n},\cdot):\mathbb{M}\to\mathbb{R}$ be the function defined by
$G_{n}(y_{n},y)\coloneqq\lambda_{n}F(y_{n},y)+d(y_{n},x_{n})[d(y,x_{n})-d(x_{n},y_{n})]\text{ for all }y\in C.$
Then $G_{n}(y_{n},y_{n})=0$. It follows from (3.6) that
$y_{n}=\operatorname*{argmin}_{y\in C}G_{n}(y_{n},y),$
which implies that $0\in\partial G_{n}({y_{n},y_{n}})$. It follows from Remark 2.10 that
$0\in\partial G_{n}({y_{n},y_{n}})$ $=$ $\partial(\lambda_{n}F(y_{n},y)+d(y_{n},x_{n})[d(y,x_{n})-d(x_{n},y_{n})])\big{|}_{y_{n}}$
$=$ $\lambda_{n}\partial F(y_{n},y_{n})+\partial(d(y_{n},x_{n})[d(y,x_{n})-d(x_{n},y_{n})])\big{|}_{y_{n}}$
$=$ $\lambda_{n}\partial F(y_{n},y_{n})+d(y_{n},x_{n})(-d^{-1}(y_{n},x_{n})\exp_{y_{n}}^{-1}x_{n})$
$=$ $\lambda_{n}\partial F(y_{n},y_{n})-\exp_{y_{n}}^{-1}x_{n}.$
Thus $(1/\lambda_{n})\exp_{y_{n}}^{-1}x_{n}\in\partial F(y_{n},y_{n})$. By the definition of $\partial F(y_{n},y_{n})$, we have
$\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x\rangle\leq\lambda_{n}(F(y_{n},x)-F(y_{n},y_{n}))=\lambda_{n}F(y_{n},x)\text{ for all }x\in\mathbb{M}.$ (3.7)
Setting $x=x_{n+1}\in C$ in (3.7), we obtain
$\lambda_{n}F(y_{n},x_{n+1})\geq\langle\exp^{-1}_{y_{n}}x_{n},\exp^{-1}_{y_{n}}x_{n+1}\rangle.$ (3.8)
Using (3.4) and (3.8), we find that
$\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x_{n+1}\rangle+\langle\exp_{x_{n+1}}^{-1}x_{n},\exp_{x_{n+1}}^{-1}y\rangle$ $\leq\lambda_{n}(F(y_{n},x_{n+1})+f(y_{n},y)-F(y_{n},x_{n+1}))$
$=\lambda_{n}f(y_{n},y)\text{ for all }y\in C.$ (3.9)
Using Remark 2.5 for $x_{n},x_{n+1},y\in\mathbb{M}$, one sees that
$2\langle\exp_{x_{n+1}}^{-1}x_{n},\exp_{x_{n+1}}^{-1}y\rangle\geq d^{2}(x_{n},x_{n+1})+d^{2}(x_{n+1},y)-d^{2}(x_{n},y)\text{ for all }y\in C.$ (3.10)
Again, using Remark 2.5 for $x_{n},y_{n},x_{n+1}\in\mathbb{M}$, one sees that
$2\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x_{n+1}\rangle\geq d^{2}(x_{n},y_{n})+d^{2}(x_{n+1},y_{n})-d^{2}(x_{n},x_{n+1}).$ (3.11)
Using (3.9), (3.10) and (3.11), we have
$d^{2}(x_{n},y_{n})+d^{2}(x_{n+1},y_{n})-d^{2}(x_{n},x_{n+1})+d^{2}(x_{n},x_{n+1})+d^{2}(x_{n+1},y)-d^{2}(x_{n},y)\leq 2\lambda_{n}f(y_{n},y),$
which implies that
$d^{2}(x_{n+1},y)$$\leq$ $d^{2}(x_{n},y)-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n})+2\lambda_{n}F(y_{n},y)\text{ for all }y\in C.$ (3.12)
Let $x^{*}\in\operatorname*{EP}(F,C)$ and put $y=x^{*}$ in (3.12). Since $y_{n}\in C$, we obtain $F(x^{*},y_{n})\geq 0$. When combined with the monotonicity of $F$, this yields $F(y_{n},x^{*})\leq 0$. Thus, using (3.12), we have
$d^{2}(x_{n+1},x^{*})\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n}),$ (3.13)
which implies that
$d(x_{n+1},x^{*})\leq d(x_{n},x^{*})\text{ for all }n\in\mathbb{N}_{0}.$ (3.14)
Since $x^{*}\in\operatorname*{EP}(F,C)$ was arbitrary, this means that $\{x_{n}\}$ is Fejér monotone with respect to $\operatorname*{EP}(F,C)$. Thus $\{x_{n}\}$ is bounded and $\lim_{n\to\infty}d(x_{n},x^{*})$ exists. By taking the limit as $n\to\infty$ in (3.13), one arrives at
$\lim_{n\to\infty}d(x_{n},y_{n})=0.$ (3.15)
Thus, the sequence $\{y_{n}\}$ is also bounded.
Theorem 3.4.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction satisfying the Assumptions (A1)-(A4). Let $\{x_{n}\}$ be the sequence generated by Algorithm 3.1 such that $\limsup_{n\to\infty}\frac{1}{\lambda_{n}}\geq\frac{1}{\lambda}$ for some $\lambda\in(0,\infty)$. Then $\{x_{n}\}$ converges to an element of $\operatorname*{EP}(F,C)$.
Proof.
In view of Lemma 3.3, the sequence $\{x_{n}\}$ is Fejér monotone with respect to the solution set $\operatorname*{EP}(F,C)$ and the sequences $\{x_{n}\}$ and $\{y_{n}\}$ are bounded. Since $\{x_{n}\}$ is bounded, it has at least one cluster point. Let $\bar{x}$ be such a cluster point, and let $\{x_{n_{k}}\}$ be a subsequence such that $x_{n_{k}}\to\bar{x}$ as $k\to\infty$. From (3.15), we have $\lim_{k\to\infty}d(x_{n_{k}},y_{n_{k}})=0$. Thus, $y_{n_{k}}\to\bar{x}$ as $k\to\infty$. Using (3.7), we see that
$F(y_{n_{k}},y)\geq\frac{1}{\lambda_{n_{k}}}\langle\exp_{y_{n_{k}}}^{-1}(x_{n_{k}}),\exp_{y_{n_{k}}}^{-1}(y)\rangle\text{ for all }y\in C.$
Taking the limit superior as $k\to\infty$ and using Assumption (A3) and Lemma 2.4, we obtain
$F(\bar{x},y)\geq\limsup_{k\to\infty}F(y_{n_{k}},y)\geq\limsup_{k\to\infty}\frac{1}{\lambda_{n_{k}}}\langle\exp_{y_{n_{k}}}^{-1}(x_{n_{k}}),\exp_{y_{n_{k}}}^{-1}(y)\rangle=0\text{ for all }y\in C,$
which implies that $\bar{x}\in\operatorname*{EP}(F,C)$. Thus, every cluster point of $\{x_{n}\}$ belongs to $\operatorname*{EP}(F,C)$. Since $\lim_{n\to\infty}d(x_{n},x^{*})$ exists for any $x^{*}\in\operatorname*{EP}(F,C)$, the entire sequence $\{x_{n}\}$ converges to $\bar{x}\in\operatorname*{EP}(F,C)$. ∎
Lemma 3.5.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction satisfying the Assumptions (A1)-(A4). Let $\{x_{n}\}$ and $\{y_{n}\}$ be the sequences generated by Algorithm 3.2. Then, for each $x^{*}\in\operatorname*{EP}(F,C)$, we have
$d^{2}(x_{n+1},x^{*})\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n}).$ (3.16)
Furthermore, $\{x_{n}\}$ is Fejér monotone with respect to the solution set $\operatorname*{EP}(F,C)$ and the sequences $\{x_{n}\}$ and $\{y_{n}\}$ are bounded.
Proof.
According to the definition of $x_{n+1}$ in Algorithm 3.1 and Remark 2.9, we have
$\lambda_{n}(F(y_{n},y)-F(y_{n},x_{n+1}))\geq\langle\exp^{-1}_{x_{n+1}}x_{n},\exp^{-1}_{x_{n+1}}y\rangle\text{ for all }y\in C.$ (3.17)
From the definition of $y_{n}$ and $J_{\lambda_{n}}^{F}(x_{n})$, it follows that
$0\leq\lambda_{n}F(y_{n},y)+(d^{2}(y,x_{n})-d^{2}(y_{n},x_{n}))\text{ for all }y\in C.$ (3.18)
Let $G_{n}(y_{n},\cdot):\mathbb{M}\to\mathbb{R}$ be the function defined by
$G_{n}(y_{n},y)=\lambda_{n}F(y_{n},y)+(d^{2}(y,x_{n})-d^{2}(y_{n},x_{n}))\text{for all }y\in C.$
Then $G_{n}(y_{n},y_{n})=0$. Using (3.18), we find that
$y_{n}=\operatorname*{argmin}_{y\in C}G_{n}(y_{n},y),$
which implies that $0\in\partial G_{n}({y_{n},y_{n}})$. It follows from Remark 2.10 that
$0\in\partial G_{n}({y_{n},y_{n}})$ $=\partial(\lambda_{n}F(y_{n},y)+(d^{2}(y,x_{n})-d^{2}(y_{n},x_{n})))\big{|}_{y_{n}}$
$=\lambda_{n}\partial F(y_{n},y_{n})+\partial(d^{2}(y,x_{n})-d^{2}(y_{n},x_{n}))\big{|}_{y_{n}}$
$=\lambda_{n}\partial F(y_{n},y_{n})-\exp_{y_{n}}^{-1}x_{n}.$
Thus $(1/\lambda_{n})\exp_{y_{n}}^{-1}x_{n}\in\partial F(y_{n},y_{n})$. By the definition of $\partial F(y_{n},y_{n})$, we have
$\frac{1}{\lambda_{n}}\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x\rangle\leq F(y_{n},x)-F(y_{n},y_{n})=F(y_{n},x)\text{ for all }x\in\mathbb{M}.$ (3.19)
Setting $x=x_{n+1}\in C$ in (3.19), we obtain
$\lambda_{n}F(y_{n},x_{n+1})\geq\langle\exp^{-1}_{y_{n}}x_{n},\exp^{-1}_{y_{n}}x_{n+1}\rangle.$ (3.20)
Using (3.17) and (3.20), we have
$\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x_{n+1}\rangle+\langle\exp_{x_{n+1}}^{-1}x_{n},\exp_{x_{n+1}}^{-1}y\rangle$ $\leq\lambda_{n}(F(y_{n},x_{n+1})+f(y_{n},y)-F(y_{n},x_{n+1}))$
$=\lambda_{n}f(y_{n},y)\text{ for all }y\in C.$ (3.21)
Using Remark 2.5 for $x_{n},x_{n+1},y\in\mathbb{M}$, one sees that
$2\langle\exp_{x_{n+1}}^{-1}x_{n},\exp_{x_{n+1}}^{-1}y\rangle\geq d^{2}(x_{n},x_{n+1})+d^{2}(x_{n+1},y)-d^{2}(x_{n},y)\text{ for all }y\in C.$ (3.22)
Again, using Remark 2.5 for $x_{n},y_{n},x_{n+1}\in\mathbb{M}$, one sees that
$2\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x_{n+1}\rangle\geq d^{2}(x_{n},y_{n})+d^{2}(x_{n+1},y_{n})-d^{2}(x_{n},x_{n+1}).$ (3.23)
Using (3.21), (3.22) and (3.23), we have
$d^{2}(x_{n},y_{n})+d^{2}(x_{n+1},y_{n})-d^{2}(x_{n},x_{n+1})+d^{2}(x_{n},x_{n+1})+d^{2}(x_{n+1},y)-d^{2}(x_{n},y)\leq 2\lambda_{n}f(y_{n},y),$
which implies that
$d^{2}(x_{n+1},y)$$\leq$ $d^{2}(x_{n},y)-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n})+2\lambda_{n}F(y_{n},y)\text{ for all }y\in C.$ (3.24)
Let $x^{*}\in\operatorname*{EP}(F,C)$ and put $y=x^{*}$ in (3.24). Since $y_{n}\in C$, we obtain $F(x^{*},y_{n})\geq 0$. When combined with the monotonicity of $F$, this shows that $F(y_{n},x^{*})\leq 0$. It now follows from (3.24) that
$d^{2}(x_{n+1},x^{*})\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n}),$ (3.25)
which implies that
$d(x_{n+1},x^{*})\leq d(x_{n},x^{*})\text{ for all }n\in\mathbb{N}_{0}.$ (3.26)
Since $x^{*}\in\operatorname*{EP}(F,C)$ was arbitrary, this means that $\{x_{n}\}$ is Fejér monotone with respect to $\operatorname*{EP}(F,C)$. Thus $\{x_{n}\}$ is bounded and $\lim_{n\to\infty}d(x_{n},x^{*})$ exists. By taking the limit as $n\to\infty$ in (3.25), one arrives at
$\lim_{n\to\infty}d(x_{n},y_{n})=0.$ (3.27)
Thus, $\{y_{n}\}$ is also bounded.
Theorem 3.6.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction satisfying the Assumptions (A1)-(A4). Let $\{x_{n}\}$ be the sequence generated by Algorithm 3.2 such that $\limsup_{n\to\infty}\frac{1}{\lambda_{n}}\geq\frac{1}{\lambda}$ for some $\lambda\in(0,\infty)$. Then $\{x_{n}\}$ converges to an element of $\operatorname*{EP}(F,C)$.
Proof.
It follows from Lemma 3.5 that $\{x_{n}\}$ is Fejér monotone with respect to the solution set $\operatorname*{EP}(F,C)$ and the sequences $\{x_{n}\}$ and $\{y_{n}\}$ are bounded. Since $\{x_{n}\}$ is bounded, it has at least one cluster point. Let $\bar{x}$ be a cluster point, and let $\{x_{n_{k}}\}$ be a subsequence such that $x_{n_{k}}\to\bar{x}$ as $k\to\infty$. Using (3.27), we have $\lim_{k\to\infty}d(x_{n_{k}},y_{n_{k}})=0$. Thus, $y_{n_{k}}\to\bar{x}$ as $k\to\infty$. Using (3.19), we have
$F(y_{n_{k}},y)\geq\frac{1}{\lambda_{n_{k}}}\langle\exp_{y_{n_{k}}}^{-1}(x_{n_{k}}),\exp_{y_{n_{k}}}^{-1}(y)\rangle\text{ for all }y\in C.$
Taking the limit as $k\to\infty$, and using Assumption (A4) and Lemma 2.4, we have
$F(\bar{x},y)\geq\limsup_{k\to\infty}f(y_{n_{k}},y)\geq\limsup_{k\to\infty}\frac{1}{\lambda_{n_{k}}}\langle\exp_{y_{n_{k}}}^{-1}(x_{n_{k}}),\exp_{y_{n_{k}}}^{-1}(y)\rangle=0\text{ for all }y\in C,$
which implies that $\bar{x}\in\operatorname*{EP}(F,C)$. Thus, every cluster point of $\{x_{n}\}$ belongs to $\operatorname*{EP}(F,C)$. Since $\lim_{n\to\infty}d(x_{n},x^{*})$ exists for any $x^{*}\in\operatorname*{EP}(F,C)$, the entire sequence $\{x_{n}\}$ converges to $\bar{x}\in\operatorname*{EP}(F,C)$. ∎
4 Error bounds and linear convergence for strongly pseudomonotone equilibrium problems
In this section we establish a global error bound and demonstrate the linear convergence of the proposed extragradient methods (Algorithms 3.1 and 3.2) for solving strongly pseudomonotone equilibrium problems.
4.1 Global error bounds
In this subsection we find the error bounds of the proposed Algorithms 3.1 and 3.2 when the involved bifunction $F$ is strongly pseudomonotone.
Theorem 4.1.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a $\beta$-strongly pseudomonotone bifunction such that $F(x,x)=0$ for all $x\in C$ and Assumptions (A2) and (A3) hold. Let $x^{*}\in\operatorname*{EP}(F,C)$ be the unique solution to the EP and let $\{x_{n}\}$ be the sequence generated by Algorithm 3.1. Then
$d(x_{n},x^{*})\leq\left(1+\frac{1}{\beta\lambda_{n}}\right)d(y_{n},x_{n})\text{ for all }n\in\mathbb{N}_{0}.$
Proof.
Using (3.7), we have
$\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x^{*}\rangle\leq\lambda_{n}F(y_{n},x^{*}).$ (4.1)
Since $x^{*}\in\operatorname*{EP}(F,C)$ and $y_{n}\in C$, one obtains $F(x^{*},y_{n})\geq 0$. Note that $F$ is $\beta$-strongly pseudomonotone. Therefore
$F(y_{n},x^{*})\leq-\beta d^{2}(y_{n},x^{*}).$ (4.2)
Using (4.1) and (4.2), we have
$\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x^{*}\rangle\leq-\beta\lambda_{n}d^{2}(y_{n},x^{*}).$ (4.3)
By the Cauchy-Schwarz inequality, we have
$\beta\lambda_{n}d^{2}(y_{n},x^{*})\leq\langle-\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x^{*}\rangle\leq\|\exp_{y_{n}}^{-1}x_{n}\|\|\exp_{y_{n}}^{-1}x^{*}\|=d(y_{n},x_{n})d(y_{n},x^{*}),$
which implies that
$d(y_{n},x^{*})\leq\frac{1}{\beta\lambda_{n}}d(y_{n},x_{n}).$
The triangle inequality implies that
$d(x_{n},x^{*})\leq d(x_{n},y_{n})+d(y_{n},x^{*})\leq d(x_{n},y_{n})+\frac{1}{\beta\lambda_{n}}d(y_{n},x_{n})=\left(1+\frac{1}{\beta\lambda_{n}}\right)d(y_{n},x_{n}).$
Theorem 4.2.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a $\beta$-strongly pseudomonotone bifunction such that $F(x,x)=0$ for all $x\in C$ and Assumptions (A2) and (A3) hold. Let $x^{*}\in\operatorname*{EP}(F,C)$ be the unique solution to the EP and let $\{x_{n}\}$ be the sequence generated by Algorithm 3.2. Then
$d(x_{n},x^{*})\leq\left(1+\frac{1}{\beta\lambda_{n}}\right)d(y_{n},x_{n})\text{ for all }n\in\mathbb{N}_{0}.$
Proof.
The proof follows similar steps to those in the proof of Theorem 4.1. ∎
4.2 Linear convergence
In this subsection we establish the $R$-linear convergence rates of the proposed Algorithms 3.1 and 3.2 when the involved bifunction $F$ is strongly pseudomonotone.
Definition 4.3.
A sequence $\{x_{n}\}$ in an Hadamard manifold $\mathbb{M}$ is said to converge $R$-linearly to $x^{*}$ with rate $\eta\in[0,1)$ if there exists a constant $c>0$ such that $d(x_{n},x^{*})\leq c\eta^{n}$ for all $n\in\mathbb{N}$.
Theorem 4.4.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a $\beta$-strongly pseudomonotone bifunction such that $F(x,x)=0$ for all $x\in C$ and Assumptions (A2) and (A3) hold. Let $\{x_{n}\}$ be the sequence generated by Algorithm 3.1 such that $\{\lambda_{n}\}\subset(\tilde{\lambda},\infty)$ and $\limsup_{n\to\infty}\frac{1}{\lambda_{n}}\geq\frac{1}{\lambda}$ for some $\lambda,\tilde{\lambda}\in(0,\infty)$. Then $\{x_{n}\}$ converges $R$-linearly to the unique solution of EP with the rate $r=\sqrt{(1-\min\{1,2\beta\tilde{\lambda}\})}$.
Proof.
Under the above assumption, the equilibrium problem $EP(F,C)$ has a unique solution, say $x^{*}$. Since $x^{*}\in\operatorname*{EP}(F,C)$ and $y_{n}\in C$, one obtains $F(x^{*},y_{n})\geq 0$, which, when combined with the $\beta$-strong pseudomonotonicity of $F$ implies that $F(y_{n},x^{*})\leq-\beta d^{2}(y_{n},x^{*})$. Letting $y=x^{*}$ in (3.12), one has
$d^{2}(x_{n+1},x^{*})$ $\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n})+2\lambda_{n}F(y_{n},x^{*})$ (4.4)
$\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n})-2\beta\lambda_{n}d^{2}(y_{n},x^{*}).$
$\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-2\beta\lambda_{n}d^{2}(y_{n},x^{*}).$
$\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-2\beta\tilde{\lambda}d^{2}(y_{n},x^{*}).$
$\leq d^{2}(x_{n},x^{*})-\min\{1,2\beta\tilde{\lambda}\}(d^{2}(x_{n},y_{n})+d^{2}(y_{n},x^{*})).$
$\leq d^{2}(x_{n},x^{*})-\min\{1,2\beta\tilde{\lambda}\}(d^{2}(x_{n},x^{*})).$
$=(1-\min\{1,2\beta\tilde{\lambda}\})d^{2}(x_{n},x^{*})=r^{2}d^{2}(x_{n},x^{*}),$ (4.5)
where $r=\sqrt{(1-\min\{1,2\beta\tilde{\lambda}\})}\in[0,1)$. Thus, we have
$d(x_{n+1},x^{*})\leq rd(x_{n},x^{*})\leq\cdots\leq r^{n+1}d(x_{0},x^{*}).$ (4.6)
That is,
$d(x_{n+1},x^{*})\leq cr^{n+1},$
where $c=d(x_{0},x^{*})$. In other words $\{x_{n}\}$ is $R$-linearly convergent with rate $r$. ∎
Theorem 4.5.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a $\beta$-strongly pseudomonotone bifunction such that $F(x,x)=0$ for all $x\in C$ and Assumptions (A2) and (A3) hold. Let $\{x_{n}\}$ be the sequence generated by Algorithm 3.2 such that $\{\lambda_{n}\}\subset(\tilde{\lambda},\infty)$ and $\limsup_{n\to\infty}\frac{1}{\lambda_{n}}\geq\frac{1}{\lambda}$ for some $\lambda,\tilde{\lambda}\in(0,\infty)$. Then $\{x_{n}\}$ converges $R$-linearly to the unique solution of EP with the rate $r=\sqrt{(1-\min\{1,2\beta\tilde{\lambda}\})}$.
Proof.
The proof follows similar steps to those in the proof of Theorem 4.4. ∎
5 Numerical Experiments
In this section we present numerical examples to illustrate the computational performance of the algorithms proposed in this paper and compare them with existing ones. All codes were written in MATLAB 2024b and executed on a MacBook Air with Intel Core i5-5350U (1.8 GHz, dual-core) processor and 8 GB RAM.
Let ++$\mathbb{R}^{N}_{++}=\{x=(x_{1},\ldots,x_{N})\in\mathbb{R}^{N}:x_{i}>0,~{}i=1,\ldots,N\}$ and let $\mathbb{M}=(\mathbb{R}^{N}_{++},\langle\cdot,\cdot\rangle)$ be a Riemannian manifold with the Riemannian metric defined by
$\langle u,v\rangle=uG(x)v^{T},\quad x\in\mathbb{R}^{N}_{++},u,v\in T_{x}\mathbb{R}^{N}_{++}=\mathbb{R}^{N},$ (5.1)
where $G(x)$ is a diagonal matrix defined by $G(x)=\text{diag}(x_{1}^{-2},\ldots,x_{N}^{-2})$. The Riemannian distance $d:\mathbb{M}\times\mathbb{M}\to[0,\infty)$ is defined by
$d(x,y)=\left(\sum_{i=1}^{N}\ln^{2}\frac{x_{i}}{y_{i}}\right)^{1/2}\text{ for all }x,y\in\mathbb{M}.$
The sectional curvature of this Riemannian manifold $\mathbb{M}$ is 00. Thus $M=(\mathbb{R}^{N}_{++},\langle\cdot,\cdot\rangle)$ is an Hadamard manifold.
Example 5.1.
Let $\mathbb{M}$ be the Riemannian manifold with the Riemannian metric defined by (5.1). Let $N=3$ and let $F:\mathbb{M}\times\mathbb{M}\to\mathbb{R}$ be defined by
$F(x,y)$ $=\left(3\ln\left(\frac{x_{1}x_{2}}{x_{3}}\right)\ln\left(\frac{y_{1}}{x_{1}}\right),3\ln\left(\frac{x_{1}x_{2}}{x_{3}}\right)\ln\left(\frac{y_{2}}{x_{2}}\right),\right.$
$\quad\left.-3\ln\left(\frac{x_{1}x_{2}}{x_{3}}\right)\ln\left(\frac{y_{3}}{x_{3}}\right)\right)\text{ for all }x,y\in\mathbb{M}.$
Note that $F$ satisfies the Assumptions (A1)-(A4) and the resolvent $J_{\lambda}^{F}:\mathbb{M}\to\mathbb{M}$ is given by
$J_{\lambda}^{F}(x)=((x_{1}x_{2}^{3\lambda}x_{3}^{3\lambda})^{\frac{1}{1+3\lambda}},(x_{1}^{3\lambda}x_{2}^{-1}x_{3}^{3\lambda})^{\frac{1}{1+3\lambda}},(x_{1}^{3\lambda}x_{2}^{3\lambda}x_{3}^{1+6\lambda})^{\frac{1}{1+3\lambda}})\text{ for all }x\in\mathbb{M}.$
We conduct a numerical experiment and apply our Algorithm 3.1, denoted by Alg. REM (Regularized Extragradient Method), to approximate the solution of the equilibrium problem (EP) (1.1), where the bifunction involved is given by $F$ as defined in Example 5.1. We choose the parameter sequence $\{\lambda_{n}\}$ to be constant, i.e., $\lambda_{n}=\lambda$ for all $n\in\mathbb{N}$. Ten different values of $\lambda$ are selected using the MATLAB command 3*linspace(0.01, 0.1, 10). Figure LABEL:Fig1 illustrates the results for two different initial points: $x_{0}=(1,2,3)$ and $x_{0}=(5,5,5)$. Box plots are also presented for different initial points by MATLAB command randi([5, 20], N, 1) with 30 independent trials. For all plots and tables, our stopping criterion is defined as $Er(n)=d(x_{n+1},x_{n})\leq 10^{-8}$.
Table 1 presents the mean iteration counts and computational times across multiple trials with varying initial points. As $\lambda$ increases, both the number of iterations and computational time decrease significantly, indicating faster convergence of the method. In particular, the best performance is observed for $\lambda=0.30$, where Alg. REM converges in approximately 31 iterations on average. The iteration plots in Figure LABEL:Fig1 further confirm this trend for the different initial points.
For each value of $\lambda$, we present box plots of the number of iterations and the computational time in Figure LABEL:Fig2 (a) and Figure LABEL:Fig2 (b), respectively. In each box plot, the central red line denotes the median, while the bottom and top edges of the box correspond to the 25th and 75th percentiles. The whiskers extend to the most extreme data points not considered outliers, and outliers are shown individually using the ‘+’ symbol. As observed in Figure LABEL:Fig2 (a), the number of iterations decreases steadily as $\lambda$ increases, indicating faster convergence for larger values of $\lambda$. A similar trend is evident in Figure LABEL:Fig2 (b), where computational time also decreases significantly with increasing $\lambda$, demonstrating the efficiency gains associated with larger regularization parameters.
$\lambda$ Mean Iter. Std Dev Mean Time (s) Std Dev 0.030 202.37 3.10 0.0031 0.0080 0.060 110.10 1.58 0.0008 0.0023 0.090 77.80 1.65 0.0008 0.0012 0.120 61.23 1.57 0.0007 0.0012 0.150 51.80 0.81 0.0003 0.0003 0.180 44.73 0.98 0.0002 0.0002 0.210 39.93 0.94 0.0001 0.0001 0.240 36.27 0.94 0.0001 0.0001 0.270 33.57 0.62 0.0001 0.0001 0.300 31.40 0.57 0.0001 0.0001
Table 1: Numerical results for Alg. REM with different values of $\lambda$ and randomly generated initial points
In the following example, we compare the numerical behavior of Algorithm 3.1 from Tan et al. [31], denoted as Alg. TQY, with that of Alg. REM.
Example 5.2.
Let $\mathbb{M}$ be the Riemannian manifold with the Riemannian metric defined by (5.1) and let $F:\mathbb{M}\times\mathbb{M}\to\mathbb{R}$ be defined by
$F(x,y)=\left(\ln x_{1}\ln\left(\frac{y_{1}}{x_{1}}\right),\ldots,\ln x_{N}\ln\left(\frac{y_{N}}{x_{N}}\right)\right)\text{ for all }x,y\in\mathbb{M}.$
Then $F$ is monotone and $F(x,x)=0$. Also note that all the conditions $(A2)$-$(A4)$ are satisfied and the resolvent $J_{\lambda}^{F}:\mathbb{M}\to\mathbb{M}$ is given by
$J_{\lambda}^{F}(x)=(x_{1}^{\frac{1}{1+\lambda}},\ldots,x_{N}^{\frac{1}{1+\lambda}})\text{ for all }x\in\mathbb{M}.$
We set the problem dimension to $N=100$ and $N=1000$ and choose initial point $x_{0}$ generated randomly using the MATLAB command randi([5, 20], N, 1). For Alg. REM, we use a constant parameter sequence $\{\lambda_{n}\}$ selected via 3*linspace(0.01, 0.1, 10). For Alg. TQY, we set $\tau_{0}$ from the same range, while fixing $\xi_{n}=0$ and $\sigma_{n}=\frac{1}{(n+5)^{2}}$ to analyze the algorithm’s sensitivity to the initial regularization parameter. The stopping criterion is defined by $Er(n)=d(x_{n+1},x_{n})\leq 10^{-8}$.
Table 2 presents the average iteration count and computational time for both algorithms across various values of $\lambda$ (or $\tau_{0}$ for Alg. TQY). The results show a consistent advantage of Alg. REM over Alg. TQY in terms of both number of iterations and computational time. As $\lambda$ increases, Alg. REM exhibits significantly fewer iterations and reduced computational time.
To further analyze the convergence behavior of both algorithms, we present iteration plots and box plots illustrating the number of iterations and computational time required to meet the stopping criterion. The iteration plots for the initial point $x_{0}=ones(N,1)$ and $\lambda=0.03$ clearly demonstrate that Alg. REM exhibits significantly faster and more stable convergence compared to Alg. TQY across all tested values of the regularization parameter, see Figure LABEL:Fig. The box plots provide additional statistical insight into the variability of performance. For Alg. REM, as $\lambda$ increases, the spread of the box plots narrows considerably, indicating reduced variability in both iteration count and computational time across multiple independent trials; see Figure LABEL:BoxItr and LABEL:BoxCT. This reflects robust and stable behavior of Alg. REM for larger parameter values. On the other hand, Alg. TQY algorithm exhibits broader boxes and more frequent outliers, suggesting higher variability and less predictable convergence. These visualizations collectively confirm that Alg. REM not only achieves faster convergence but also maintains more consistent performance, particularly when larger regularization parameters are employed.
Dimension (100) $\lambda=0.03$ $\lambda=0.09$ $\lambda=0.15$ $\lambda=0.21$ $\lambda=0.30$ Algorithms REM TQY REM TQY REM TQY REM TQY REM TQY Av. Itr.(n) 421 622 145 617 89 611 66 610 48 606 Av. Time(s) 0.014670 0.032896 0.004084 0.022525 0.001953 0.019044 0.001570 0.019195 0.001044 0.018917 Dimension (1000) $\lambda=0.03$ $\lambda=0.09$ $\lambda=0.15$ $\lambda=0.21$ $\lambda=0.30$ Algorithms REM TQY REM TQY REM TQY REM TQY REM TQY Av. Itr.(n) 460 683 158 677 98 679 72 678 52 676 Av. Time(s) 0.057001 0.117077 0.016142 0.112079 0.010240 0.111515 0.008309 0.114897 0.005299 0.114566
Table 2: Numerical results for Alg. REM, and Alg. TQY for $N=100,1000$ and randomly generated initial points
Acknowledgment
No funding or external support was received for this research.
Disclosure statement
Competing of Interest: The authors declare that they have no competing interests.
Data Availability: Availability of data and materials not applicable.