### v2

Abstract
Employing two distinct types of regularization terms, we propose two regularized extragradient methods for solving equilibrium problems on Hadamard manifolds. The sequences generated by these extragradient algorithms converge to a solution of the equilibrium problem without requiring the Lipschitz continuity of the bifunction or imposing additional conditions on the parameters. We establish convergence results for both algorithms and derive global error bounds along with $R$-linear convergence rates in cases where the bifunction is strongly pseudomonotone. Finally, we present numerical experiments to demonstrate the effectiveness of our methods.
Keywords: Hadamard manifold, Equilibrium Problem, Busemann function, Extragradient method, Linear convergence
1 Introduction
The equilibrium problem (EP) has been extensively studied and remains a highly active area of research. One of the key motivations behind its investigation is its broad applicability, since numerous mathematical problems can be formulated as equilibrium problems. These include optimization problems, Nash equilibrium problems, complementarity problems, fixed point problems, and variational inequality problems. A comprehensive discussion of EP and its developments can be found in the works of Blum and Oettli [Blum1994], Binachi and Schaible [Bianchi1996], as well as in references therein.
The study of the proximal point method on Hadamard manifolds has gained considerable attention, with various researchers exploring its application to specific cases of the equilibrium problem (EP); see, for instance, the works of Ferreira and Oliveira [Ferreira2002], Li et al. [Li2009JLM], and Tang et al. [Tang2013]. More recently, there has been a natural shift toward extending classical Euclidean concepts and methodologies to the Riemannian setting. Notable contributions in this direction include those by Ledyaev and Zhu [Ledyaev2007], Ferreira et al. [Ferreira2005], Li et al. [Li2011SJO, Li2011SVVA], among others. A significant advantage of adapting these techniques to Riemannian manifolds is the ability to reformulate certain nonconvex problems as convex ones by selecting an appropriate Riemannian metric. Relevant discussions on this topic can be found in the papers by Rapcsák [Rapscak], Cruz Neto et al. [Neto2002], Bento and Melo [Bento2012], and by Colao et al. [Colao2012].
Note that Hadamard manifolds generally do not have a linear structure, which indicates that properties, techniques as well as algorithms in linear spaces cannot be used in Hadamard manifolds. Therefore, it is valuable and interesting to extend algorithms for solving equilibrium problems from linear spaces to Hadamard manifolds.
Let $\mathbb{M}$ be an Hadamard manifold. Given $C\subset\mathbb{M}$, a nonempty, closed and convex set, and a bifunction $F:C\times C\to\mathbb{R}$ satisfying the property $F(x,x)=0$ for all $x\in C$, the equilibrium problem in the Riemanninan context (denoted by $\operatorname*{EP}(F,C)$) is formulated as follows:
$\text{Find }x^{*}\in C\text{ such that }F(x^{*},y)\geq 0\text{ for all }y\in C.$ (1)
In recent years various extragradient-based methods have been developed in order to solve equilibrium problem (1) on Hadamard manifolds. Colao et al. [Colao2012] established the existence of solutions to such problems under suitable conditions on the bifunction. Building upon their work, Cruz Neto et al. [Neto2016OL] proposed an extragradient algorithm with a fixed step size and proved its global convergence. Inspired by the works of Cruz Neto et al. [Neto2016OL] and Hieu [Hieu2018AA, Hieu2019], Khammahawong et al. [Khammahawong2020] introduced two extragradient methods employing decreasing and non-summable step sizes, extending Hieu’s results from Hilbert space to the Hadamard setting. For strongly pseudomonotone bifunctions, they demonstrated convergence of the generated sequences under mild assumptions.
Further extending this line of research, Ansari et al. [Ansari2020] proposed two adaptive explicit extragradient algorithms for solving EP (1) on Hadamard manifolds. Their methods utilize a simple, information-based adaptive step size rule, eliminating the need for prior knowledge of the Lipschitz constant. Convergence was established under pseudomonotonicity, with linear convergence shown for strongly pseudomonotone bifunctions.
It is worth noting that the methods in [Neto2016OL, Khammahawong2020, Ansari2020] require solving a strongly convex optimization problem twice per iteration. To reduce the computational cost, Chen et al. [Chen2021] proposed a modified golden ratio algorithm that requires only one evaluation per iteration. Numerical experiments have shown that their method outperforms those of [Neto2016OL, Khammahawong2020] in terms of efficiency and computational cost.
In 2022, Babu et al. [Babu2022] developed several extragradient algorithms incorporating an Armijo-type step size rule for solving equilibrium problems on Hadamard manifolds. More recently, Oyewole and Reich [Oyewole2023inertial] introduced an inertial subgradient extragradient method with a self-adaptive step size. However, these methods tend to be computationally intensive, as the associated strongly convex optimization problem may need to be solved multiple times at each iteration.
In 2022, Bento et al. [Bento2022AOR] introduced a regularization of the proximal point algorithm in the context of Hadamard manifolds using a new regularization term which has a clear interpretation in a recent variational rationality approach to human behavior. Later, Bento et al. [Bento2022] introduced a new resolvent for the bifunction in terms of Busemann functions. An advantage of this new proposal is that, in addition to being a natural extension of its counterpart in the linear setting introduced by Combettes and Hirstoaga [Hirstoaga], the new term that performs the regularization is a convex function in general Hadamard manifolds. Bento et al. [Bento2024] have recently considered the resolvent via Busemann functions introduced by Bento et al. [Bento2022] and presented a proximal point method for solving equilibrium problems on Hadamard manifolds.
Motivated by the works of Bento et al. [Bento2022AOR], Bento et al. [Bento2022] and Bento et al. [Bento2024], we propose two regularized extragradient methods incorporating distinct regularization terms for solving equilibrium problems on Hadamard manifolds. The proposed algorithms operate without any step size restrictions, do not require a line search procedure, and do not assume Lipschitz continuity of the bifunction, thus making them suitable for solving equilibrium problems involving monotone bifunctions.
2 Preliminaries
For the basic definitions and results regarding the geometry of Riemannian manifolds, we refer to [Docarmo, Sakai, Udriste]. Let $\mathbb{M}$ be an $m$-dimensional differentiable manifold and let $x\in\mathbb{M}$. The set of all tangent vectors at the point $x$, denoted by $T_{x}\mathbb{M}$, is called the tangent space of $\mathbb{M}$ at $x$. It forms a real vector space of dimension $m$. The tangent bundle of $\mathbb{M}$ is defined as the disjoint union $T\mathbb{M}=\bigcup_{x\in\mathbb{M}}T_{x}\mathbb{M}$.
We suppose that $\mathbb{M}$ is equipped with a Riemannian metric, that is, a smoothly varying family of inner products on the tangent spaces $T_{x}\mathbb{M}$. The metric at $x\in\mathbb{M}$ is denoted $\langle\cdot,\cdot\rangle_{x}:T_{x}\mathbb{M}\times T_{x}\mathbb{M}\to\mathbb{R}$. A differentiable manifold $\mathbb{M}$ with a Riemannian metric $\langle\cdot,\cdot\rangle$ is said to be a Riemannian manifold. The norm corresponding to the inner product $\langle\cdot,\cdot\rangle_{x}$ on $T_{x}\mathbb{M}$ is denoted by $\|\cdot\|_{x}$.
Let $x,y\in\mathbb{M}$ and let $\gamma:[a,b]\to\mathbb{M}$ be a piecewise smooth curve joining $x$ to $y$. Then the length of the curve $\gamma$ is defined by $L(\gamma):=\int_{a}^{b}\|\gamma^{\prime}(t)\|dt$, where $\gamma^{\prime}(t)\in T_{\gamma(t)}\mathbb{M}$ is a tangent vector at the point $\gamma(t)\in\mathbb{M}$. The minimal length of all such curves joining $x$ to $y$ is called the Riemannian distance. It induces the original topology on $\mathbb{M}$ and is denoted by $d(x,y)$. In other words, $d(x,y):=\inf\{L(\gamma):\gamma\in C_{xy}\}$. where $C_{xy}$ denotes the set of all continuously differentiable curves $\gamma:[0,1]\to\mathbb{M}$ such that $\gamma(0)=x$ and $\gamma(1)=y$.
A single-valued vector field on a differentiable manifold $\mathbb{M}$ is a smooth mapping $A\colon\mathbb{M}\to T\mathbb{M}$ such that, for each $x\in\mathbb{M}$, there exists a tangent vector $A(x)\in T_{x}\mathbb{M}$. Let $\triangle$ be the Levi-Civita connection associated with the Riemannian manifold $M$. A vector field $A$ along $\gamma$ is said to be parallel if $\triangle_{\gamma^{\prime}}A=0$. If $\gamma^{\prime}$ is parallel along $\gamma$, that is, $\triangle_{\gamma^{\prime}}\gamma^{\prime}=0$, then $\gamma$ is said to be a geodesic. In this case, $\|\gamma^{\prime}\|$ is a constant. Moreover, if $\|\gamma^{\prime}\|=1$, then $\gamma$ is called a normalized geodesic. A geodesic joining $x$ to $y$ in the Riemannian manifold $\mathbb{M}$ is said to be a minimal geodesic if its length is equal to $d(x,y)$. We denote the geodesic joining $x$ and $y$ by $\gamma(x,y;\cdot)$, that is, $\gamma(x,y;\cdot):[a,b]\to\mathbb{M}$ is such that, for $a,b\in\mathbb{R}$, we have $\gamma(x,y;a)=x$ and $\gamma(x,y;b)=y$.
The Riemannian metric induces a mapping $f\mapsto\operatorname*{grad}f$ that associates to each differentiable function $f\colon\mathbb{M}\to\mathbb{R}$ a unique vector field $\operatorname*{grad}f\colon\mathbb{M}\to T\mathbb{M}$, called the gradient of $f$, defined by $\langle\operatorname*{grad}f(x),X(x)\rangle=df_{x}(X(x))$ for all vector fields $X\colon\mathbb{M}\to T\mathbb{M}$ and all $x\in\mathbb{M}$. Here, $df_{x}$ denotes the differential of $f$ at the point $x\in\mathbb{M}$.
A Riemannian manifold $\mathbb{M}$ is said to be complete if for any $x\in\mathbb{M}$, each geodesic emanating from $x$ is defined for all $t\in\mathbb{R}$. By the Hopf-Rinow theorem [Sakai, Theorem 1.1], if $\mathbb{M}$ is complete, then each pair of points in $\mathbb{M}$ can be joined by a minimal geodesic. Moreover, $(\mathbb{M},d)$ is a complete metric space. If $\mathbb{M}$ is a complete Riemannian manifold, then the exponential map $\exp_{x}\colon T_{x}\mathbb{M}\to\mathbb{M}$ at $x\in\mathbb{M}$ is defined by $\exp_{x}v=\gamma_{x,v}(1)\text{ for all }v\in T_{x}\mathbb{M}$, where $\gamma_{x,v}\colon\mathbb{R}\to\mathbb{M}$ is the unique geodesic starting from $x$ with velocity $v$, that is, $\gamma_{x,v}(0)=x$ and $\gamma^{\prime}_{x,v}(0)=v$. It is known that $\exp_{x}(tv)=\gamma_{x,v}(t)$ for each real number $t$ and $\exp_{x}(\mathbf{0})=\gamma_{x,\mathbf{0}}(0)=x$, where $\mathbf{0}$ is the zero tangent vector.
Definition 1 ([Martin2]).
A complete simply connected Riemannian manifold of nonpositive sectional curvature is called an Hadamard manifold.
Definition 2 ([Martin2]).
A nonempty subset $C$ of an Hadamard manifold $\mathbb{M}$ is said to be geodesically convex if, for any two points $x$ and $y$ in $C$, the geodesic joining $x$ to $y$ is contained in $C$, that is, if $\gamma(x,y;\cdot)\colon[a,b]\to\mathbb{M}$ is the geodesic such that $x=\gamma(x,y;a)$ and $y=\gamma(x,y;b)$, then $\gamma((1-t)a+tb)\in C$ for all $t\in[0,1]$.
Definition 3 ([Martin2]).
Let $\mathbb{M}$ be an Hadamard manifold. A function $f\colon\mathbb{M}\to\mathbb{R}$ is said to be geodesically convex if, for any geodesic $\gamma(x,y;\cdot)\colon[a,b]\to\mathbb{M}$, the composite function $f\circ\gamma(x,y;\cdot)\colon[a,b]\to\mathbb{R}$ is convex, that is,
$f\circ\gamma(x,y;(1-t)a+tb)\leq(1-t)(f\circ\gamma(x,y;a))+t(f\circ\gamma(x,y;b))$
for any $a,b\in\mathbb{R}$, $x,y\in\mathbb{M}$ and $t\in[a,b]$.
Lemma 1 ([Martin2]).
Let $\{x_{n}\}$ be a sequence in an Hadamard manifold $\mathbb{M}$ such that $x_{n}\to x\in\mathbb{M}$. Then, for any $y\in\mathbb{M}$, we have
[…]
Remark 1 ([Martin2]).
Let $\mathbb{M}$ be an Hadamard manifold and let $x,y,z\in\mathbb{M}$. Then
$d^{2}(x,y)+d^{2}(y,z)-d^{2}(z,x)\leq 2\langle\exp_{y}^{-1}x,\exp_{y}^{-1}z\rangle.$ (2)
Definition 4 ([Udriste]).
Let $\mathbb{M}$ be an Hadamard manifold and let $C$ be a nonempty, geodesically convex subset of $\mathbb{M}$. Let $g\colon C\to\mathbb{R}$ be a real-valued function. Then the subdifferential $\partial g\colon\mathbb{M}\to T\mathbb{M}$ of $g$ at $x$ is defined by
$\partial g(x)\coloneqq\left\{v\in T_{x}\mathbb{M}:\langle v,\exp_{x}^{-1}y\rangle\leq g(y)-g(x)\text{ for all }y\in C\right\}.$
Definition 5 ([Jost1995]).
Let $C$ be a nonempty subset of an Hadamard manifold $\mathbb{M}$ and let $f\colon C\to(-\infty,\infty]$ be a proper, geodesically convex and lower semicontinuous function. The proximal operator of $f$ is defined by
$\operatorname*{prox}_{\lambda f}(x)\coloneq\operatorname*{argmin}_{y\in C}\left(f(y)+\frac{1}{2\lambda}d^{2}(y,x)\right)\text{ for all }x\in C\text{ and for any }\lambda\in(0,\infty).$
Lemma 2 ([Ferreira2002], Theorem 5.1).
Let $\mathbb{M}$ be an Hadamard manifold and let $f:\mathbb{M}\to\mathbb{R}$ be a geodesically convex function. Let $\{x_{n}\}$ be the sequence generated by the proximal point algorithm, that is,
$x_{n+1}=\operatorname*{argmin}_{y\in\mathbb{M}}\left\{f(y)+\frac{1}{2\lambda_{n}}d^{2}(x_{n},y)\right\}\text{ for all }n\in\mathbb{N}_{0}$
with the initial point $x_{0}\in\mathbb{M}$ and $\lambda_{n}\in(0,\infty)$. Then the sequence $\{x_{n}\}$ is well defined and characterized by $\frac{1}{\lambda_{n}}\exp_{x_{n+1}}^{-1}x_{n}\in\partial f(x_{n+1}).$
Remark 2.
In view of [Ferreira2002, Lemma 4.2], $\operatorname*{prox}_{\lambda f}(x)$ is single-valued. By the definition of $\partial f(x_{n+1})$ and Lemma 2, we have
$\frac{1}{\lambda_{n}}\langle\exp_{x_{n+1}}^{-1}x_{n},\exp_{x_{n+1}}^{-1}x\rangle\leq f(x)-f(x_{n+1})\text{ for all }x\in\mathbb{M}.$
Remark 3.
Note that $\operatorname*{grad}d(y,x_{n})=-\frac{\exp_{y}^{-1}x_{n}}{d(y,x_{n})}$ provided $y\neq x_{n}$.
Definition 6 ([Bento2022AOR, Tan2024]).
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$. A bifunction $F:C\times C\to\mathbb{R}$ is said to be
  1. (i)
    monotone if $F(x,y)+F(y,x)\leq 0$ for all $x,y\in C$;
  2. (ii)
    pseudomonotone if $F(x,y)\geq 0$ implies $F(y,x)\leq 0$ for all $x,y\in C$;
  3. (iii)
    $\alpha$-strongly monotone if there exists a constant $\alpha>0$ such that $F(x,y)+F(y,x)\leq-\alpha d^{2}(x,y)$ for all $x,y\in C$;
  4. (iv)
    $\beta$-strongly pseudomonotone if there exists a constant $\beta>0$ such that $F(x,y)\geq 0$ implies $F(y,x)\leq-\beta d^{2}(x,y)$ for all $x,y\in C$.
The Busemann functions, originally introduced by Herbert Busemann to formulate the parallel axiom in certain classes of metric spaces [Busemann1955], have since found applications far beyond their original purpose; see, for example, [Busemann1993, Li1987, Shiohama, Sormani] and the references therein.
Definition 7 ([Busemann1993]).
Let $\mathbb{M}$ be an Hadamard manifold and let $\gamma_{z,x}:[0,\infty)\to\mathbb{M}$ be a geodesic ray parametrized by arc length, starting from $z$ and passing through $x$. The Busemann function associated with $\gamma_{z,x}$ is defined by
$b_{\gamma_{z,x}}(y)\coloneqq\lim_{t\to+\infty}\left[d(y,\gamma_{z,x}(t))-t\right],\quad\text{for all }y\in\mathbb{M}.$
Recently, Busemann functions have been studied in the context of optimization; see, [Bento2022, Bento2023]. In this direction, Bento et al. [Bento2022] introduced the following resolvent associated with a bifunction, which is defined using a Busemann function.
Definition 8 ([Bento2022]).
Let $C$ be a nonempty, closed, and geodesically convex subset of an Hadamard manifold $\mathbb{M}$, and let $F:C\times C\to\mathbb{R}$ be a bifunction. The Busemann resolvent $J_{\lambda}^{F}:C\to\mathbb{M}$ associated with $F$ is defined by
$J_{\lambda}^{F}(x)\coloneqq\{z\in C:\lambda F(z,y)+d(z,x)b_{\gamma_{z,x}}(y)\geq 0\text{ for all }y\in C\},$ (3)
for all $x\in C$ and any $\lambda\in(0,\infty)$.
By [Ballmann, Lemma 3.4], the Busemann function $b_{\gamma_{x,z}}(\cdot)$ is Lipschitz continuous with Lipschitz constant equal to $1$, geodesically convex, and satisfies $\|\operatorname{grad}b_{\gamma_{x,z}}(\cdot)\|=1$. Moreover, as noted in [Bridson, page 273], in the Euclidean setting one has
$d(x,z)\,b_{\gamma_{x,z}}(y)=\langle x-z,y-x\rangle.$
This identity shows that the additional term appearing in (3) is a natural extension of the regularization term introduced in [Hirstoaga].
On the other hand, it is known that any linear affine function on the Poincaré plane must be constant; see [Udriste, Theorem 2.2, p. 299]. This fact reflects the absence of nontrivial affine functions in negatively curved spaces. In line with [Bento2023, Theorem 1], this result extends to general Hadamard manifolds with negative sectional curvature. Consequently, Busemann functions provide a meaningful substitute for affine functions in linear settings. As a result, the resolvent defined in (3), which incorporates a Busemann function, serves as a natural and well-adapted analogue of the classical linear resolvent in the Riemannian framework.
Proposition 3.
Let $\mathbb{M}$ be an Hadamard manifold and let $x,z\in\mathbb{M}$. Then
$d(x,z)b_{\gamma_{z,x}}(y)\geq-\langle\exp_{z}^{-1}x,\exp_{z}^{-1}y\rangle\text{ for all }y\in\mathbb{M}.$
Proof.
Since $\gamma_{z,x}$ is the geodesic ray parameterized by arc length, we have
$\gamma_{z,x}(t))=\exp_{z}(t/d(x,z))\exp_{z}^{-1}x\text{ for all }t\in[0,\infty).$
Using (2), we get
$d^{2}(y,\gamma_{z,x}(t))$ $\geq d^{2}(z,y)+d^{2}(\gamma_{z,x}(t),z)-2\langle\exp_{z}^{-1}\gamma_{z,x}(t),\exp_{z}^{-1}y\rangle$
$=d^{2}(z,y)+\frac{t^{2}}{d^{2}(x,z)}d^{2}(z,x)-2\frac{t}{d(x,z)}\langle\exp_{z}^{-1}x,\exp_{z}^{-1}y\rangle$
$=d^{2}(z,y)-\frac{2t}{d(x,z)}\langle\exp_{z}^{-1}x,\exp_{z}^{-1}y\rangle-t^{2}.$ (4)
Using the same argument as in [Bridson, Example 8.24 (3), Page 274] and (4), we have
$b_{\gamma_{z,x}}(y)$ $=\lim_{t\to+\infty}\left[d(y,\gamma_{z,x}(t))-t\right]=\lim_{t\to+\infty}\frac{1}{2t}\left(d^{2}(y,\gamma_{z,x}(t))-t^{2}\right)$
$\geq\lim_{t\to+\infty}\frac{1}{2t}\left(d^{2}(z,y)-\frac{2t}{d(x,z)}\langle\exp_{z}^{-1}x,\exp_{z}^{-1}y\rangle\right)$
$=-\frac{1}{d(x,z)}\langle\exp_{z}^{-1}x,\exp_{z}^{-1}y\rangle.$
Therefore $d(x,z)b_{\gamma_{z,x}}(y)\geq-\langle\exp_{z}^{-1}x,\exp_{z}^{-1}y\rangle$. ∎
Remark 4.
Let $\mathbb{M}$ be an Hadamard manifold with zero sectional curvature. Then
  1. (i)
    Let $x,z\in\mathbb{M}$. Using Proposition 3, we have
    $d(x,z)\,b_{\gamma_{z,x}}(y)=-\langle\exp_{z}^{-1}x,\;\exp_{z}^{-1}y\rangle\text{ for all }y\in\mathbb{M}.$
  2. (ii)
    From [Colao2012, Theorem 4.7(ii)], it follows that $J_{\lambda}^{F}$ is firmly nonexpansive.
Assumption 4.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction. We consider the following assumptions:
  1. (A1)
    $F$ is monotone and $F(x,x)=0$ for all $x\in C$;
  2. (A2)
    $F(x,\cdot)$ is geodesically convex and subdifferentiable on $C$ for each $x\in C$;
  3. (A3)
    $F(\cdot,y)$ is upper semicontinuous on $C$ for each $y\in C$, that is, $\limsup_{n\to\infty}F(x_{n},y)\leq F(x,y)$ for each $y\in C$ and each $\{x_{n}\}\subset C$ with $x_{n}\to x$;
  4. (A4)
    Given a fixed $z_{0}\in\mathbb{M}$, consider a sequence $\{z_{n}\}\subset C$ such that $d(z_{n},z_{0})\to\infty$ as $n\to\infty$. Then there exist $x^{*}\in C$ and $n_{0}\in\mathbb{N}$ such that
    $F(z_{n},x^{*})\leq 0\text{ for all }n\geq n_{0}.$
Let $\lambda\in(0,\infty)$ and $x\in C$ and let $F:C\times C\to\mathbb{R}$ be a bifunction. For each $x\in C$, consider the set
$K_{x}\coloneqq\{z\in C:\lambda F(z,y)+(d^{2}(y,z)-d^{2}(x,z))\geq 0\}.$
Then, using [Bento2022, Theorem 4.1], [Bento2024, Theorem 1.1], and [Bento2022AOR, Corollary 1], we arrive at the following result:
Theorem 5.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction such that Assumptions (A1) - (A4) hold. Let $\lambda\in(0,\infty)$. Then
$\operatorname*{EP}(F,C)\neq\emptyset$,
$J_{\lambda}^{F}(x)\neq\emptyset$ for all $x\in\mathbb{M}$,
$J_{\lambda}^{F}$ is single-valued,
the set $K_{x}\neq\emptyset$ for each $x\in C$,
the set $K_{x}$ is a singleton for each $x\in C$.
Theorem 6 ([Tan2024], Theorem 4.1).
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a $\beta$-strongly monotone bifunction such that Assumptions (A2) and (A3) hold. Then equilibrium problem (1) has a unique solution.
Definition 9 ([Ferreira2002]).
Let $C$ be a nonempty subset of a complete metric space $X$. A sequence $\{x_{n}\}\subset X$ is called Fejér monotone with respect to $C$ if
$d(x_{n+1},y)\leq d(x_{n},y)\text{ for all }y\in C\text{ and for all }n\in\mathbb{N}_{0}.$
Lemma 7 ([Ferreira2002]).
Let $C$ be a nonempty subset of a complete metric space $X$. If $\{x_{n}\}\subset X$ is Fejér monotone with respect to $C$, then $\{x_{n}\}$ is bounded. In addition, if $\{x_{n}\}$ has a cluster point $x$ in $C$, then $\{x_{n}\}$ converges to $x$.
3 Regularized extragradient methods for equilibrium problems
In this section we present two regularized extragradient methods for solving equilibrium problem (1) on Hadamard manifolds.
Algorithm 8.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction.
Initialization: Choose $x_{0}\in C$ and $\lambda_{0}\in(0,\infty)$.
Iterative Steps: Given the current iterate $x_{n}$, compute the next iterate as follows:
$\begin{cases}y_{n}=\left\{x\in C:\lambda_{n}F(x,y)+d(x,x_{n})b_{\gamma_{x,x_{n}}}(y)\geq 0\text{ for all }y\in C\right\}=J_{\lambda_{n}}^{F}(x_{n}),\\ x_{n+1}=\operatorname*{argmin}_{y\in C}\left\{F(y_{n},y)+\frac{1}{2\lambda_{n}}d^{2}(x_{n},y)\right\}=\operatorname*{prox}_{\lambda_{n}F(y_{n},\cdot)}(x_{n})\text{ for all }n\in\mathbb{N}_{0},\end{cases}$ (5)
where $\{\lambda_{n}\}$ is a sequence in $(0,\infty)$ such that $\limsup_{n\to\infty}\frac{1}{\lambda_{n}}\geq\frac{1}{\lambda}$ for some $\lambda\in(0,\infty)$.
Algorithm 9.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction.
Initialization: Choose $x_{0}\in C$ and $\lambda_{0}\in(0,\infty)$.
Iterative Steps: Given the current iterate $x_{n}$, compute the next iterate as follows:
$\begin{cases}y_{n}=\left\{x\in C:\lambda_{n}F(x,y)+(d^{2}(y,x_{n})-d^{2}(x,x_{n}))\geq 0\text{ for all }y\in C\right\},\\ x_{n+1}=\operatorname*{argmin}_{y\in C}\left\{F(y_{n},y)+\frac{1}{2\lambda_{n}}d^{2}(x_{n},y)\right\}=\operatorname*{prox}_{\lambda_{n}F(y_{n},\cdot)(x_{n})}\text{ for all }n\in\mathbb{N}_{0},\end{cases}$ (6)
where $\{\lambda_{n}\}$ is a sequence in $(0,\infty)$ such that $\limsup_{n\to\infty}\frac{1}{\lambda_{n}}\geq\frac{1}{\lambda}$ for some $\lambda\in(0,\infty)$.
The well-definedness of Algorithms 8 and 9 follows from Theorem 5. Next, we prove an important lemma that is essential for the convergence analysis of Algorithm 8.
Lemma 10.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction satisfying Assumptions (A1)-(A4). Let $\{x_{n}\}$ and $\{y_{n}\}$ be sequences generated by Algorithm 8. Then, for each $x^{*}\in\operatorname*{EP}(F,C)$, we have
$d^{2}(x_{n+1},x^{*})\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n}).$ (7)
Furthermore, $\{x_{n}\}$ is Fejér monotone with respect to $\operatorname*{EP}(F,C)$ and the sequences $\{x_{n}\}$ and $\{y_{n}\}$ are bounded.
Proof.
Using the definition of $x_{n+1}$ in Algorithm 8 and Remark 2, we have
$\lambda_{n}(F(y_{n},y)-F(y_{n},x_{n+1}))\geq\langle\exp^{-1}_{x_{n+1}}x_{n},\exp^{-1}_{x_{n+1}}y\rangle\text{ for all }y\in C.$ (8)
Using the definition of $y_{n}$ and $J_{\lambda_{n}}^{F}(x_{n})$, we have
$0\leq\lambda_{n}F(y_{n},y)+d(y_{n},x_{n})b_{\gamma_{y_{n},x_{n}}}(y)\text{ for all }y\in C.$ (9)
Since $b_{\gamma_{y_{n},x_{n}}}(y)\leq d(y,x_{n})-d(x_{n},y_{n})$ for all $y\in C$, by (9), we have
$0\leq\lambda_{n}F(y_{n},y)+d(y_{n},x_{n})[d(y,x_{n})-d(x_{n},y_{n})]\text{ for all }y\in C.$ (10)
Let $G_{n}(y_{n},\cdot):\mathbb{M}\to\mathbb{R}$ be the function defined by
$G_{n}(y_{n},y)\coloneqq\lambda_{n}F(y_{n},y)+d(y_{n},x_{n})[d(y,x_{n})-d(x_{n},y_{n})]\text{ for all }y\in C.$
Then $G_{n}(y_{n},y_{n})=0$. It follows from (10) that
$y_{n}=\operatorname*{argmin}_{y\in C}G_{n}(y_{n},y),$
which implies that $0\in\partial G_{n}({y_{n},y_{n}})$. It follows from Remark 3 that
$0\in\partial G_{n}({y_{n},y_{n}})$ $=$ $\partial(\lambda_{n}F(y_{n},y)+d(y_{n},x_{n})[d(y,x_{n})-d(x_{n},y_{n})])\big|_{y_{n}}$
$=$ $\lambda_{n}\partial F(y_{n},y_{n})+\partial(d(y_{n},x_{n})[d(y,x_{n})-d(x_{n},y_{n})])\big|_{y_{n}}$
$=$ $\lambda_{n}\partial F(y_{n},y_{n})+d(y_{n},x_{n})(-d^{-1}(y_{n},x_{n})\exp_{y_{n}}^{-1}x_{n})$
$=$ $\lambda_{n}\partial F(y_{n},y_{n})-\exp_{y_{n}}^{-1}x_{n}.$
Thus $(1/\lambda_{n})\exp_{y_{n}}^{-1}x_{n}\in\partial F(y_{n},y_{n})$. By the definition of $\partial F(y_{n},y_{n})$, we have
$\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x\rangle\leq\lambda_{n}(F(y_{n},x)-F(y_{n},y_{n}))=\lambda_{n}F(y_{n},x)\text{ for all }x\in\mathbb{M}.$ (11)
Setting $x=x_{n+1}\in C$ in (11), we obtain
$\lambda_{n}F(y_{n},x_{n+1})\geq\langle\exp^{-1}_{y_{n}}x_{n},\exp^{-1}_{y_{n}}x_{n+1}\rangle.$ (12)
Using (8) and (12), we find that
$\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x_{n+1}\rangle+\langle\exp_{x_{n+1}}^{-1}x_{n},\exp_{x_{n+1}}^{-1}y\rangle$ $\leq\lambda_{n}(F(y_{n},x_{n+1})+F(y_{n},y)-F(y_{n},x_{n+1}))$
$=\lambda_{n}F(y_{n},y)\text{ for all }y\in C.$ (13)
Using (2) for $x_{n},x_{n+1},y\in\mathbb{M}$, one sees that
$2\langle\exp_{x_{n+1}}^{-1}x_{n},\exp_{x_{n+1}}^{-1}y\rangle\geq d^{2}(x_{n},x_{n+1})+d^{2}(x_{n+1},y)-d^{2}(x_{n},y)\text{ for all }y\in C.$ (14)
Again, using (2) for $x_{n},y_{n},x_{n+1}\in\mathbb{M}$, one sees that
$2\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x_{n+1}\rangle\geq d^{2}(x_{n},y_{n})+d^{2}(x_{n+1},y_{n})-d^{2}(x_{n},x_{n+1}).$ (15)
Using (13), (14) and (15), we have
$d^{2}(x_{n},y_{n})+d^{2}(x_{n+1},y_{n})-d^{2}(x_{n},x_{n+1})+d^{2}(x_{n},x_{n+1})+d^{2}(x_{n+1},y)-d^{2}(x_{n},y)\leq 2\lambda_{n}F(y_{n},y),$
which implies that
$d^{2}(x_{n+1},y)$ $\leq$ $d^{2}(x_{n},y)-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n})+2\lambda_{n}F(y_{n},y)\text{ for all }y\in C.$ (16)
Let $x^{*}\in\operatorname*{EP}(F,C)$ and put $y=x^{*}$ in (16). Since $y_{n}\in C$, we obtain $F(x^{*},y_{n})\geq 0$. When combined with the monotonicity of $F$, this yields $F(y_{n},x^{*})\leq 0$. Thus, using (16), we have
$d^{2}(x_{n+1},x^{*})\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n}),$ (17)
which implies that
$d(x_{n+1},x^{*})\leq d(x_{n},x^{*})\text{ for all }n\in\mathbb{N}_{0}.$ (18)
Since $x^{*}\in\operatorname*{EP}(F,C)$ is arbitrary, this means that $\{x_{n}\}$ is Fejér monotone with respect to $\operatorname*{EP}(F,C)$. By Lemma 7, the sequence $\{x_{n}\}$ is bounded, and $\lim_{n\to\infty}d(x_{n},x^{*})$ exists. By taking the limit as $n\to\infty$ in (17), one arrives at
$\lim_{n\to\infty}d(x_{n},y_{n})=0.$ (19)
Thus, the sequence $\{y_{n}\}$ is also bounded. ∎
Theorem 11.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction satisfying Assumptions (A1)-(A4). Let $\{x_{n}\}$ be the sequence generated by Algorithm 8 where $\limsup_{n\to\infty}\frac{1}{\lambda_{n}}\geq\frac{1}{\lambda}$ for some $\lambda\in(0,\infty)$. Then $\{x_{n}\}$ converges to an element of $\operatorname*{EP}(F,C)$.
Proof.
In view of Lemma 10, the sequence $\{x_{n}\}$ is Fejér monotone with respect to the solution set $\operatorname*{EP}(F,C)$ and the sequences $\{x_{n}\}$ and $\{y_{n}\}$ are bounded. Since $\{x_{n}\}$ is bounded, it has at least one cluster point. Let $\bar{x}$ be such a cluster point, and let $\{x_{n_{k}}\}$ be a subsequence such that $x_{n_{k}}\to\bar{x}$ as $k\to\infty$. From (19), we have $\lim_{k\to\infty}d(x_{n_{k}},y_{n_{k}})=0$. Thus, $y_{n_{k}}\to\bar{x}$ as $k\to\infty$. Using (11), we see that
$F(y_{n_{k}},y)\geq\frac{1}{\lambda_{n_{k}}}\langle\exp_{y_{n_{k}}}^{-1}(x_{n_{k}}),\exp_{y_{n_{k}}}^{-1}(y)\rangle\text{ for all }y\in C.$
Taking the limit superior as $k\to\infty$ and using Assumption (A3) and Lemma 1, we obtain
[…]
which implies that $\bar{x}\in\operatorname*{EP}(F,C)$. Thus, every cluster point of $\{x_{n}\}$ belongs to $\operatorname*{EP}(F,C)$. Since $\lim_{n\to\infty}d(x_{n},x^{*})$ exists for any $x^{*}\in\operatorname*{EP}(F,C)$, the entire sequence $\{x_{n}\}$ converges to $\bar{x}\in\operatorname*{EP}(F,C)$. ∎
We now turn to the convergence analysis of Algorithm 9, for which the following lemma is essential.
Lemma 12.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction satisfying Assumptions (A1)-(A4). Let $\{x_{n}\}$ and $\{y_{n}\}$ be the sequences generated by Algorithm 9. Then, for each $x^{*}\in\operatorname*{EP}(F,C)$, we have
$d^{2}(x_{n+1},x^{*})\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n}).$ (20)
Furthermore, $\{x_{n}\}$ is Fejér monotone with respect to the solution set $\operatorname*{EP}(F,C)$ and the sequences $\{x_{n}\}$ and $\{y_{n}\}$ are bounded.
Proof.
According to the definition of $x_{n+1}$ in Algorithm 8 and Remark 2, we have
$\lambda_{n}(F(y_{n},y)-F(y_{n},x_{n+1}))\geq\langle\exp^{-1}_{x_{n+1}}x_{n},\exp^{-1}_{x_{n+1}}y\rangle\text{ for all }y\in C.$ (21)
From the definition of $y_{n}$ and $J_{\lambda_{n}}^{F}(x_{n})$, it follows that
$0\leq\lambda_{n}F(y_{n},y)+(d^{2}(y,x_{n})-d^{2}(y_{n},x_{n}))\text{ for all }y\in C.$ (22)
Let $G_{n}(y_{n},\cdot):\mathbb{M}\to\mathbb{R}$ be the function defined by
$G_{n}(y_{n},y)=\lambda_{n}F(y_{n},y)+(d^{2}(y,x_{n})-d^{2}(y_{n},x_{n}))\text{ for all }y\in C.$
Then $G_{n}(y_{n},y_{n})=0$. Using (22), we find that
$y_{n}=\operatorname*{argmin}_{y\in C}G_{n}(y_{n},y),$
which implies that $0\in\partial G_{n}({y_{n},y_{n}})$. It follows from Remark 3 that
$0\in\partial G_{n}({y_{n},y_{n}})$ $=\partial(\lambda_{n}F(y_{n},y)+(d^{2}(y,x_{n})-d^{2}(y_{n},x_{n})))\big|_{y_{n}}$
$=\lambda_{n}\partial F(y_{n},y_{n})+\partial(d^{2}(y,x_{n})-d^{2}(y_{n},x_{n}))\big|_{y_{n}}$
$=\lambda_{n}\partial F(y_{n},y_{n})-\exp_{y_{n}}^{-1}x_{n}.$
Thus $(1/\lambda_{n})\exp_{y_{n}}^{-1}x_{n}\in\partial F(y_{n},y_{n})$. By the definition of $\partial F(y_{n},y_{n})$, we have
$\frac{1}{\lambda_{n}}\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x\rangle\leq F(y_{n},x)-F(y_{n},y_{n})=F(y_{n},x)\text{ for all }x\in\mathbb{M}.$ (23)
Setting $x=x_{n+1}\in C$ in (23), we obtain
$\lambda_{n}F(y_{n},x_{n+1})\geq\langle\exp^{-1}_{y_{n}}x_{n},\exp^{-1}_{y_{n}}x_{n+1}\rangle.$ (24)
Using (21) and (24), we have
$\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x_{n+1}\rangle+\langle\exp_{x_{n+1}}^{-1}x_{n},\exp_{x_{n+1}}^{-1}y\rangle$ $\leq\lambda_{n}(F(y_{n},x_{n+1})+f(y_{n},y)-F(y_{n},x_{n+1}))$
$=\lambda_{n}f(y_{n},y)\text{ for all }y\in C.$ (25)
Using (2) for $x_{n},x_{n+1},y\in\mathbb{M}$, one sees that
$2\langle\exp_{x_{n+1}}^{-1}x_{n},\exp_{x_{n+1}}^{-1}y\rangle\geq d^{2}(x_{n},x_{n+1})+d^{2}(x_{n+1},y)-d^{2}(x_{n},y)\text{ for all }y\in C.$ (26)
Again, using (2) for $x_{n},y_{n},x_{n+1}\in\mathbb{M}$, one sees that
$2\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x_{n+1}\rangle\geq d^{2}(x_{n},y_{n})+d^{2}(x_{n+1},y_{n})-d^{2}(x_{n},x_{n+1}).$ (27)
Using (25), (26) and (27), we have
$d^{2}(x_{n},y_{n})+d^{2}(x_{n+1},y_{n})-d^{2}(x_{n},x_{n+1})+d^{2}(x_{n},x_{n+1})+d^{2}(x_{n+1},y)-d^{2}(x_{n},y)\leq 2\lambda_{n}f(y_{n},y),$
which implies that
$d^{2}(x_{n+1},y)$ $\leq$ $d^{2}(x_{n},y)-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n})+2\lambda_{n}F(y_{n},y)\text{ for all }y\in C.$ (28)
Let $x^{*}\in\operatorname*{EP}(F,C)$ and put $y=x^{*}$ in (28). Since $y_{n}\in C$, we obtain $F(x^{*},y_{n})\geq 0$. When combined with the monotonicity of $F$, this shows that $F(y_{n},x^{*})\leq 0$. It now follows from (28) that
$d^{2}(x_{n+1},x^{*})\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n}),$ (29)
which implies that
$d(x_{n+1},x^{*})\leq d(x_{n},x^{*})\text{ for all }n\in\mathbb{N}_{0}.$ (30)
Since $x^{*}\in\operatorname*{EP}(F,C)$ was arbitrary, this means that $\{x_{n}\}$ is Fejér monotone with respect to $\operatorname*{EP}(F,C)$. By Lemma 7, the sequence $\{x_{n}\}$ is bounded, and $\lim_{n\to\infty}d(x_{n},x^{*})$ exists. By taking the limit as $n\to\infty$ in (29), one arrives at
$\lim_{n\to\infty}d(x_{n},y_{n})=0.$ (31)
Thus, $\{y_{n}\}$ is also bounded. ∎
Theorem 13.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a bifunction satisfying the Assumptions (A1)-(A4). Let $\{x_{n}\}$ be the sequence generated by Algorithm 9 where $\limsup_{n\to\infty}\frac{1}{\lambda_{n}}\geq\frac{1}{\lambda}$ for some $\lambda\in(0,\infty)$. Then $\{x_{n}\}$ converges to an element of $\operatorname*{EP}(F,C)$.
Proof.
It follows from Lemma 12 that $\{x_{n}\}$ is Fejér monotone with respect to the solution set $\operatorname*{EP}(F,C)$ and the sequences $\{x_{n}\}$ and $\{y_{n}\}$ are bounded. Since $\{x_{n}\}$ is bounded, it has at least one cluster point. Let $\bar{x}$ be a cluster point, and let $\{x_{n_{k}}\}$ be a subsequence such that $x_{n_{k}}\to\bar{x}$ as $k\to\infty$. Using (31), we have $\lim_{k\to\infty}d(x_{n_{k}},y_{n_{k}})=0$. Thus, $y_{n_{k}}\to\bar{x}$ as $k\to\infty$. Using (23), we have
$F(y_{n_{k}},y)\geq\frac{1}{\lambda_{n_{k}}}\langle\exp_{y_{n_{k}}}^{-1}(x_{n_{k}}),\exp_{y_{n_{k}}}^{-1}(y)\rangle\text{ for all }y\in C.$
Taking the limit as $k\to\infty$, and using Assumption (A4) and Lemma 1, we have
[…]
which implies that $\bar{x}\in\operatorname*{EP}(F,C)$. Thus, every cluster point of $\{x_{n}\}$ belongs to $\operatorname*{EP}(F,C)$. Since $\lim_{n\to\infty}d(x_{n},x^{*})$ exists for any $x^{*}\in\operatorname*{EP}(F,C)$, the entire sequence $\{x_{n}\}$ converges to $\bar{x}\in\operatorname*{EP}(F,C)$. ∎
4 Error bounds and linear convergence for strongly pseudomonotone equilibrium problems
In this section we establish a global error bound and demonstrate the linear convergence of the proposed extragradient methods (Algorithms 8 and 9) for solving strongly pseudomonotone equilibrium problems.
4.1 Global error bounds
In this subsection we find error bounds for the proposed Algorithms 8 and 9 when the involved bifunction $F$ is strongly pseudomonotone.
Theorem 14.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a $\beta$-strongly pseudomonotone bifunction such that $F(x,x)=0$ for all $x\in C$ and Assumptions (A2) and (A3) hold. Let $x^{*}\in\operatorname*{EP}(F,C)$ be the unique solution to the EP and let $\{x_{n}\}$ be the sequence generated by Algorithm 8. Then
$d(x_{n},x^{*})\leq\left(1+\frac{1}{\beta\lambda_{n}}\right)d(y_{n},x_{n})\text{ for all }n\in\mathbb{N}_{0}.$
Proof.
Using (11), we have
$\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x^{*}\rangle\leq\lambda_{n}F(y_{n},x^{*}).$ (32)
Since $x^{*}\in\operatorname*{EP}(F,C)$ and $y_{n}\in C$, one obtains $F(x^{*},y_{n})\geq 0$. Note that $F$ is $\beta$-strongly pseudomonotone. Therefore
$F(y_{n},x^{*})\leq-\beta d^{2}(y_{n},x^{*}).$ (33)
Using (32) and (33), we have
$\langle\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x^{*}\rangle\leq-\beta\lambda_{n}d^{2}(y_{n},x^{*}).$ (34)
By the Cauchy-Schwarz inequality, we have
$\beta\lambda_{n}d^{2}(y_{n},x^{*})\leq\langle-\exp_{y_{n}}^{-1}x_{n},\exp_{y_{n}}^{-1}x^{*}\rangle\leq\|\exp_{y_{n}}^{-1}x_{n}\|\|\exp_{y_{n}}^{-1}x^{*}\|=d(y_{n},x_{n})d(y_{n},x^{*}),$
which implies that
$d(y_{n},x^{*})\leq\frac{1}{\beta\lambda_{n}}d(y_{n},x_{n}).$
The triangle inequality implies that
$d(x_{n},x^{*})\leq d(x_{n},y_{n})+d(y_{n},x^{*})\leq d(x_{n},y_{n})+\frac{1}{\beta\lambda_{n}}d(y_{n},x_{n})=\left(1+\frac{1}{\beta\lambda_{n}}\right)d(y_{n},x_{n}).$
Theorem 15.
Let $C$ be a nonempty, closed and geodesically convex subset of an Hadamard manifold $\mathbb{M}$ and let $F:C\times C\to\mathbb{R}$ be a $\beta$-strongly pseudomonotone bifunction such that $F(x,x)=0$ for all $x\in C$ and Assumptions (A2) and (A3) hold. Let $x^{*}\in\operatorname*{EP}(F,C)$ be the unique solution to the EP and let $\{x_{n}\}$ be the sequence generated by Algorithm 9. Then
$d(x_{n},x^{*})\leq\left(1+\frac{1}{\beta\lambda_{n}}\right)d(y_{n},x_{n})\text{ for all }n\in\mathbb{N}_{0}.$
Proof.
The proof follows similar steps to those in the proof of Theorem 14. ∎
4.2 Linear convergence
In this subsection we establish the $R$-linear convergence rates of the proposed Algorithms 8 and 9 when the involved bifunction $F$ is strongly pseudomonotone.
Definition 10.
A sequence $\{x_{n}\}$ in an Hadamard manifold $\mathbb{M}$ is said to converge $R$-linearly to $x^{*}$ with rate $\eta\in[0,1)$ if there exists a constant $c>0$ such that $d(x_{n},x^{*})\leq c\eta^{n}$ for all $n\in\mathbb{N}$.
Theorem 16.
[…]
Proof.
Under the above assumption, the equilibrium problem $EP(F,C)$ has a unique solution, say $x^{*}$. Since $x^{*}\in\operatorname*{EP}(F,C)$ and $y_{n}\in C$, one obtains $F(x^{*},y_{n})\geq 0$, which, when combined with the $\beta$-strong pseudomonotonicity of $F$, implies that $F(y_{n},x^{*})\leq-\beta d^{2}(y_{n},x^{*})$. Letting $y=x^{*}$ in (16), one has
$d^{2}(x_{n+1},x^{*})$ $\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n})+2\lambda_{n}F(y_{n},x^{*})$ (35)
$\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-d^{2}(x_{n+1},y_{n})-2\beta\lambda_{n}d^{2}(y_{n},x^{*}).$
$\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-2\beta\lambda_{n}d^{2}(y_{n},x^{*}).$
$\leq d^{2}(x_{n},x^{*})-d^{2}(x_{n},y_{n})-2\beta\tilde{\lambda}d^{2}(y_{n},x^{*}).$
$\leq d^{2}(x_{n},x^{*})-\min\{1,2\beta\tilde{\lambda}\}(d^{2}(x_{n},y_{n})+d^{2}(y_{n},x^{*})).$
$\leq d^{2}(x_{n},x^{*})-\min\{1,2\beta\tilde{\lambda}\}(d^{2}(x_{n},x^{*})).$
$=(1-\min\{1,2\beta\tilde{\lambda}\})d^{2}(x_{n},x^{*})=r^{2}d^{2}(x_{n},x^{*}),$ (36)
where $r=\sqrt{(1-\min\{1,2\beta\tilde{\lambda}\})}\in[0,1)$. Thus, we have
$d(x_{n+1},x^{*})\leq rd(x_{n},x^{*})\leq\cdots\leq r^{n+1}d(x_{0},x^{*}).$ (37)
That is,
$d(x_{n+1},x^{*})\leq cr^{n+1},$
where $c=d(x_{0},x^{*})$. In other words, $\{x_{n}\}$ is $R$-linearly convergent with rate $r$. ∎
Theorem 17.
[…]
Proof.
The proof follows similar steps to those in the proof of Theorem 16. ∎
5 Numerical Experiments
In this section we present numerical examples to illustrate the computational performance of our proposed algorithms and compare them with existing ones. All codes were written in MATLAB 2024b and executed on a MacBook Air with Intel Core i5-5350U (1.8 GHz, dual-core) processor and 8 GB RAM.
Let $\mathbb{R}^{N}_{++}=\{x=(x_{1},\ldots,x_{N})\in\mathbb{R}^{N}:x_{i}>0,~i=1,\ldots,N\}$, and let $\mathbb{M}=(\mathbb{R}^{N}_{++},\langle\cdot,\cdot\rangle)$ be a Riemannian manifold with the Riemannian metric defined by
$\langle u,v\rangle=uG(x)v^{T},\quad x\in\mathbb{R}^{N}_{++},u,v\in T_{x}\mathbb{R}^{N}_{++}=\mathbb{R}^{N},$ (38)
where $G(x)$ is a diagonal matrix defined by $G(x)=\text{diag}(x_{1}^{-2},\ldots,x_{N}^{-2})$. The Riemannian distance $d:\mathbb{M}\times\mathbb{M}\to[0,\infty)$ is defined by
$d(x,y)=\left(\sum_{i=1}^{N}\ln^{2}\frac{x_{i}}{y_{i}}\right)^{1/2}\text{ for all }x,y\in\mathbb{M}.$
The sectional curvature of this Riemannian manifold $\mathbb{M}$ is 0. Thus $M=(\mathbb{R}^{N}_{++},\langle\cdot,\cdot\rangle)$ is an Hadamard manifold; see [Shikher2024NA].
Let $x,z\in\mathbb{M}$. Then the Busemann function $b_{\gamma_{z,x}}\colon\mathbb{M}\to\mathbb{R}$ is given by
$b_{\gamma_{z,x}}(y)=\frac{\sum_{i=1}^{N}\ln\left(\frac{x_{i}}{z_{i}}\right)\ln\left(\frac{z_{i}}{y_{i}}\right)}{\left(\sum_{i=1}^{N}\ln^{2}\frac{x_{i}}{z_{i}}\right)^{1/2}}\text{ for all }y\in\mathbb{M}.$
Example 1.
Let $\mathbb{M}$ be the Riemannian manifold with the Riemannian metric defined by (38). Let $N=3$ and let $F:\mathbb{M}\times\mathbb{M}\to\mathbb{R}$ be defined by
$F(x,y)$ $=\left(3\ln\left(\frac{x_{1}x_{2}}{x_{3}}\right)\ln\left(\frac{y_{1}}{x_{1}}\right)+3\ln\left(\frac{x_{1}x_{2}}{x_{3}}\right)\ln\left(\frac{y_{2}}{x_{2}}\right)\right$.
$\quad\left.-3\ln\left(\frac{x_{1}x_{2}}{x_{3}}\right)\ln\left(\frac{y_{3}}{x_{3}}\right)\right)\text{ for all }x,y\in\mathbb{M}.$
Note that $F$ satisfies Assumptions (A1)-(A4) and the resolvent $J_{\lambda}^{F}:\mathbb{M}\to\mathbb{M}$ is given by
$J_{\lambda}^{F}(x)=((x_{1}x_{2}^{3\lambda}x_{3}^{3\lambda})^{\frac{1}{1+3\lambda}},(x_{1}^{3\lambda}x_{2}^{-1}x_{3}^{3\lambda})^{\frac{1}{1+3\lambda}},(x_{1}^{3\lambda}x_{2}^{3\lambda}x_{3}^{1+6\lambda})^{\frac{1}{1+3\lambda}})\text{ for all }x\in\mathbb{M}.$
For each $x\in\mathbb{M}$, the set $K_{x}$ is given by
$K_{x}=\{((x_{1}x_{2}^{3\lambda/2}x_{3}^{3\lambda/2})^{\frac{1}{1+3\lambda/2}},(x_{1}^{3\lambda/2}x_{2}^{-1}x_{3}^{3\lambda/2})^{\frac{1}{1+3\lambda/2}},(x_{1}^{3\lambda/2}x_{2}^{3\lambda/2}x_{3}^{1+6\lambda/2})^{\frac{1}{1+3\lambda/2}})\}.$
We conduct a numerical experiment and apply our Algorithm 8, denoted by Alg. REMB, and Algorithm 9, denoted by Alg. REMD, to find the solution of equilibrium problem (1), where the bifunction is given by $F$ as defined in Example 1. We choose the parameter sequence $\{\lambda_{n}\}$ to be constant, that is, $\lambda_{n}=\lambda$ for all $n\in\mathbb{N}$. Ten different values of $\lambda$ are selected using the MATLAB command 3*linspace(0.01, 0.1, 10). Figures 1 and 3 illustrate the results for two different initial points: $x_{0}=(1,2,3)$ and $x_{0}=(5,5,5)$. Box plots are also presented for different initial points by the MATLAB command randi([5, 20], N, 1) with 30 independent trials, see Figure 2 and 4. For all plots and tables, our stopping criterion is defined as $Er(n)=d(x_{n+1},x_{n})\leq 10^{-16}$.
Tables 1 and 2 present the mean iteration counts and computational times across multiple trials with varying initial points for Alg. REMB and Alg. REMD, respectively. As $\lambda$ increases, both the number of iterations and computational time decrease significantly. In particular, the best performance is observed for $\lambda=0.30$, where Algorithm REMB converges in approximately 31 iterations on average, while Algorithm REMD requires about 22 iterations, indicating faster convergence of REMD compared to REMB. The iteration plots in Figures 1 and 3 further confirm this trend for the different initial points.
For each value of $\lambda$, we present box plots of the number of iterations and the computational time in Figures 2 and Figure 4, respectively. In each box plot, the central red line denotes the median, while the bottom and top edges of the box correspond to the 25th and 75th percentiles. The whiskers extend to the most extreme data points not considered outliers, and outliers are shown individually using the ‘+’ symbol. As observed in Figures 2 (a) and 4 (a), the number of iterations decreases steadily as $\lambda$ increases, indicating faster convergence for larger values of $\lambda$. A similar trend is evident in Figures 2 (b) and 4 (b), where computational time also decreases significantly with increasing $\lambda$, demonstrating the efficiency gains associated with larger regularization parameters.
lambda Mean Iterations Std Dev Mean Time (s) Std Dev
0.03 202 3.46 0.000686 0.000299
0.06 110 1.65 0.000334 0.000063
0.09 78 1.40 0.000295 0.000536
0.12 62 0.99 0.000180 0.000068
0.15 51 1.22 0.000079 0.000004
0.18 45 1.02 0.000069 0.000002
0.21 40 0.73 0.000064 0.000004
0.24 36 0.50 0.000057 0.000003
0.27 34 0.56 0.000057 0.000010
0.30 31 0.52 0.000049 0.000003
Table 1: Numerical results for Alg. REMB with different values of $\lambda$ and randomly generated initial points
lambda Mean Iterations Std Dev Mean Time (s) Std Dev
0.03 204 1.68 0.000758 0.000398
0.06 106 0.91 0.000280 0.000078
0.09 72 0.61 0.000118 0.000017
0.12 55 0.32 0.000131 0.000028
0.15 44 0.51 0.000079 0.000019
0.18 37 0.41 0.000063 0.000008
0.21 32 0.00 0.000049 0.000008
0.24 28 0.00 0.000043 0.000003
0.27 25 0.00 0.000039 0.000002
0.30 22 0.41 0.000035 0.000001
Table 2: Numerical results for Alg. REMD with different values of $\lambda$ and randomly generated initial points
(a) $x_{0}=(1,2,3)$
(b) $x_{0}=(5,5,5)$
Figure 1: Convergence behavior of Alg. REMB for different values of $\lambda$ with initial points $x_{0}=(1,2,3)$ and $x_{0}=(5,5,5)$
(a) Box plot of iterations
(b) Box plot of computational time
Figure 2: Box plots of Alg. REMB for different values of $\lambda$ and random initial points
(a) $x_{0}=(1,2,3)$
(b) $x_{0}=(5,5,5)$
Figure 3: Convergence behavior of Alg. REMD for different values of $\lambda$ with initial points $x_{0}=(1,2,3)$ and $x_{0}=(5,5,5)$
(a) Box plot of iterations
(b) Box plot of computational time
Figure 4: Box plots of Alg. REMD for different values of $\lambda$ and random initial points
In the following example we compare the numerical behavior of Algorithm 3.1 from Tan et al. [Tan2024], denoted as Alg. TQY, with that of Alg. REMB and Alg. REMD.
Example 2.
Let $\mathbb{M}$ be the Riemannian manifold with the Riemannian metric defined by (38) and let $F:\mathbb{M}\times\mathbb{M}\to\mathbb{R}$ be defined by
$F(x,y)=\left(\ln x_{1}\ln\left(\frac{y_{1}}{x_{1}}\right)+\cdots+\ln x_{N}\ln\left(\frac{y_{N}}{x_{N}}\right)\right)\text{ for all }x,y\in\mathbb{M}.$
Then $F$ is monotone and $F(x,x)=0$. Also note that all the conditions (A2)(A2)-$(A4)$ are satisfied and the resolvent $J_{\lambda}^{F}:\mathbb{M}\to\mathbb{M}$ is given by
$J_{\lambda}^{F}(x)=(x_{1}^{\frac{1}{1+\lambda}},\ldots,x_{N}^{\frac{1}{1+\lambda}})\text{ for all }x\in\mathbb{M}.$
For each $x\in\mathbb{M}$, the set $K_{x}$ is given by
$K_{x}=\{(x_{1}^{\frac{1}{1+\lambda/2}},\ldots,x_{N}^{\frac{1}{1+\lambda/2}})\}.$
We set the problem dimension to $N=100$ and $N=1000$ and choose initial point $x_{0}$ generated randomly using the MATLAB command randi([5, 20], N, 1). For Alg. REMB and Alg. REMD, we use a constant parameter sequence $\{\lambda_{n}\}$ selected via 3*linspace(0.01, 0.1, 10). For Alg. TQY, we set $\tau_{0}$ from the same range, while fixing $\xi_{n}=0$ and $\sigma_{n}=\frac{1}{(n+5)^{2}}$, to analyze the algorithm’s sensitivity to the initial regularization parameter. The stopping criterion is defined by $Er(n)=d(x_{n+1},x_{n})\leq 10^{-8}$.
Table 3 presents the average iteration count and computational time for all algorithms across various values of $\lambda$ (or $\tau_{0}$ for Alg. TQY). The results show a consistent advantage of Alg. REMD over Alg. REMB and Alg. TQY in terms of both number of iterations and computational time. As $\lambda$ increases, Alg. REMB exhibits significantly fewer iterations and reduced computational time compared to Alg. REMB and Alg. TQY.
To further analyze the convergence behavior of algorithms, we present iteration plots and box plots illustrating the number of iterations and computational time required to meet the stopping criterion. The iteration plots for the different values of $\lambda$ and dimension $N=100$ and $N=1000$ clearly demonstrate that Alg. REMD and Alg. REMB exhibit significantly faster and more stable convergence compared to Alg. TQY across all tested values of the regularization parameter, see Figures 5 and 6. The box plots provide additional statistical insight into the variability of performance. For Alg. REMB and Alg. REMD, as $\lambda$ increases, the spread of the box plots narrows considerably, indicating reduced variability in both iteration count and computational time across multiple independent trials; see Figures 7 and 8. This reflects robust and stable behavior of Alg. REMB and Alg. REMD for larger parameter values. On the other hand, Alg. TQY algorithm exhibits broader boxes and more frequent outliers, suggesting higher variability and less predictable convergence. These visualizations collectively confirm that Alg. REMB and Alg. REMD not only achieve faster convergence but also maintain more consistent performance, particularly when larger regularization parameters are employed.
Dimension (100) $\lambda=0.06$ $\lambda=0.18$ $\lambda=0.30$ Algorithms TQY REMB REMD TQY REMB REMD TQY REMB REMD Av. Itr.(n) 593 206 585 73 67 581 46 40 Av. Time(s) 0.22368 0.044521 0.039749 0.022525 0.001953 0.019044 0.001570 0.019195 0.001044 Dimension (1000) $\lambda=0.06$ $\lambda=0.18$ $\lambda=0.30$ Algorithms TQY REMB REMD TQY REMB REMD TQY REMB REMD Av. Itr.(n) 666 226 219 660 80 73 654 51 44 Av. Time(s) 0.043811 0.011202 0.013328 0.043704 0.0045189 0.0050687 0.041514 0.0074997 0.0090923
Table 3: Numerical results for Alg. REM, and Alg. TQY for $N=100,1000$ and randomly generated initial points
(a) $\lambda=0.06$
(b) $\lambda=0.18$
(c) $\lambda=0.3$
Figure 5: Iteration plot for dimension 100100 with initial point $x_{0}=ones(N,1)$ and different $\lambda$
(a) $\lambda=0.06$
(b) $\lambda=0.18$
(c) $\lambda=0.3$
Figure 6: Iteration plot for dimension 10001000 with initial point $x_{0}=ones(N,1)$ and different $\lambda$
(a) TQY
(b) REMB
(c) REMD
Figure 7: Iteration Box Plot
(a) TQY
(b) REMB
(c) REMD
Figure 8: Computational time Box Plot
6 Conclusion
In this work we introduced two regularized extragradient methods for solving equilibrium problems on Hadamard manifolds, employing regularization terms based on the Busemann function and a distance-type metric function. We established convergence of the proposed algorithms without assuming the Lipschitz continuity of the bifunction or imposing additional restrictions on the parameters. In addition, we derived global error bounds and proved that the generated sequences converge at an $R$-linear rate when the bifunction is strongly pseudomonotone. Moreover, we provided an approximation result for the Busemann function on Hadamard manifolds in terms of the exponential map. These theoretical findings, together with our numerical experiments, demonstrate the efficiency and robustness of the proposed methods.
Acknowledgment
No funding or external support was received for this research.
Disclosure statement
Competing of Interest: The authors declare that they have no competing interests.
Data Availability: Availability of data and materials not applicable.